### v1

# PALM: Pushing Adaptive Learning Rate Mechanisms for Continual Test-Time Adaptation

Abstract
Real-world vision models in dynamic environments face rapid shifts in domain distributions, leading to decreased recognition performance. Continual test-time adaptation (CTTA) directly adjusts a pre-trained source discriminative model to these changing domains using test data. A highly effective CTTA method involves applying layer-wise adaptive learning rates, and selectively adapting pre-trained layers. However, it suffers from the poor estimation of domain shift and the inaccuracies arising from the pseudo-labels. In this work, we aim to overcome these limitations by identifying layers through the quantification of model prediction uncertainty without relying on pseudo-labels. We utilize the magnitude of gradients as a metric, calculated by backpropagating the KL divergence between the softmax output and a uniform distribution, to select layers for further adaptation. Subsequently, for the parameters exclusively belonging to these selected layers, with the remaining ones frozen, we evaluate their sensitivity in order to approximate the domain shift, followed by adjusting their learning rates accordingly. Overall, this approach leads to a more robust and stable optimization than prior approaches. We conduct extensive image classification experiments on CIFAR-10C, CIFAR-100C, and ImageNet-C and demonstrate the efficacy of our method against standard benchmarks and prior methods.
1 Introduction
In a real-world setting, machine-perception systems [1] function in a rapidly changing environment. In such contexts, pre-trained vision models are susceptible to performance degradation caused by potential distribution shifts. For instance, while these models may have been trained on clean images, images captured by sensors could be corrupted by various weather conditions.
In the realm of addressing domain shifts due to image corruptions, test-time adaptation (TTA) has been gaining a lot of traction lately [6, 44]. Here, a pre-trained source model, with no access to its training (source) data due to privacy, is adapted to unlabeled test data through online training. One strategy in TTA involves mitigating domain shifts by adjusting the source model parameters using pseudo labels and minimizing the entropy of the model predictions [44]. While this approach effectively adapts the model to a single type of corruption, it faces two constraints when continually applied across a sequence of test tasks featuring different corruptions. Firstly, there is the issue of error accumulation: since pseudo labels are generated for unlabeled test data, errors accumulate over time. Secondly, catastrophic forgetting occurs, resulting in a long-term loss of the model’s pre-trained knowledge due to continuous adaptations and large parameter updates across tasks.
Continual test-time adaptation (CTTA) has been recently introduced to tackle these challenges. In [32, 8, 44, 37], stable optimization is achieved by updating the batch normalization (BN) parameters or re-formulating its statistics. CoTTA [45] includes full model updates, with a teacher-student network setup, to handle the above issues. To further improve the performance of CoTTA, EcoTTA [41] introduces the attachment of "meta networks", i.e., a BN layer and a convolution block, to the frozen source model for efficient adaptation.
More recently, the authors in [21] demonstrated that based on the type of distributional shift of the test data, a certain subset of layers of a pre-trained model can be "surgically" or heuristically chosen to fine-tune. Inspired by this finding, Layer-wise Auto-Weighting (LAW) [33] was introduced recently. LAW utilizes the Fisher Information Matrix (FIM) [36] to estimate domain shifts, automatically weighting trainable parameters and adjusting their learning rates accordingly. While LAW achieves state-of-the-art results for CTTA, it still encounters two significant limitations: 1) It presupposes that layer-wise importance can be accurately estimated via FIM with initially generated pseudo-labels. Nevertheless, it is widely recognized that pseudo-labels are inherently noisy and unreliable. 2) Relying on directly accumulating such approximated importance from noisy labels to calculate domain shifts can lead to inaccuracies.
In this paper, we present a novel method called PALM to push the limit of adaptive learning rate methods for CTTA. Our approach begins by identifying a subset of layers that significantly contribute to the uncertainty of model predictions. Rather than relying on pseudo-labels, we determine layer-wise importance by leveraging information from the gradient space, specifically by calculating the KL divergence between the model output and a uniform distribution. This approach eliminates dependence on pseudo-labels and enables a more accurate estimation of layer-wise importance. Subsequently, the learning rates of the parameters of these layers are adjusted based on the degree of domain shift observed in the test data. We utilize parameter sensitivity as an indicator of the degree of distribution shift and employ weighted moving averages to aggregate sensitivity from previous data batches. Compared to existing adaptive learning rate methods, our proposed approach offers a better estimation of domain shift for adjusting learning rates.
To summarise, our contributions are as follows:
  • We present PALM, a novel approach to continual test-time adaptation. Specifically, we demonstrate improved schemes for adaptive learning rate methods for CTTA.
  • Considering the prediction uncertainty of the model at test-time, we design an approach to automatically select layers that would require further adaptation and push the learning rates of parameters of these layers for model updates, by estimating the domain shift based on parameter sensitivity.
  • Our method delivers promising results in both continual and gradual TTA across three standard benchmark datasets. In addition, we conduct and demonstrate ablation experiments on the different components of our work.
2 Related Works
Test-time adaptation (TTA): The overarching target of TTA is to adapt source models to unlabelled test data in real-time in a source-free online training fashion [16, 42, 44]. TENT [44] optimizes the Batch Normalisation (BN) parameters by minimizing the Shannon entropy [40] of the pseudo labels. In [37], the statistics of the BN parameters are replaced by the ones of the input corrupted data. Adacontrast [3] combines contrastive learning and utilizes the memory module to update the entire model’s parameters. While previous TTA methods showed significant improvements, they were demonstrated for a single domain shift at a time [4, 26, 37, 44] and did not consider continual distribution shifts.
Continual TTA (CTTA): CTTA poses an adaptation challenge wherein the model must continually adjust to evolving target data, accommodating a sequence of dynamic domain shifts. The first work, CoTTA [45], considers addressing this challenging task by setting up a teacher-student network, with full model updates. The teacher model is updated using an exponential moving average (EMA) to deal with forgetting and accumulation of errors with time. EcoTTA [41] introduces "meta networks" i.e. a BN layer and a convolution block are attached to the frozen source model. At every stage, an $L_{1}$ regularisation loss between the outputs of the meta network and that of the source model is computed to handle catastrophic forgetting.
Efficient fine-tuning approaches: Freezing model parameters to learn information from downstream data has witnessed a good body of work [39, 27, 51, 22, 7, 9]. Depending upon the distribution shift of the target task, in [21], the authors demonstrated that heuristically selecting a subset of convolutional blocks for training while freezing the remaining blocks works better than full fine-tuning.
Adaptive-learning rate methods: Regarding the studies on the importance of each layer in training, [47, 49] concluded that deeper layers extract task-specific features, as opposed to the shallow layers. Based on this finding, AutoLR [35] design an adapted learning rate (LR) scheme based on the weight variations between layers. Despite attempts to set different initial learning rates for each layer, these rates remain mostly constant throughout training. In the context of CTTA, LAW [33] used the FIM [36] to calculate the domain shift and automatically adjust the learning rates of parameters. However, to obtain the layer importance, the pseudo-labels are used to compute the FIM, which would be inaccurate. In addition, a direct accumulation of FIMs across domains would lead to error accumulation. Furthermore, the LR "weight" of each layer is upper-bounded by the base LR that is used, which ignores the relative "sensitivity" of parameters towards adaptation of the current batch and domain.
Parameter sensitivity: In the literature about neural network pruning, understanding the sensitivity of parameters towards training has been studied well [19, 28, 43, 29, 23]. The overall idea is that parameters with lower sensitivity need to be adapted more with an increased learning rate, and vice-versa. [23], in addition, introduces a temporal variation of sensitivity, due to uncertainty, as another factor during the training of models for natural language understanding and machine translation tasks.
3 Background
3.1 Problem Setting
The objective of continual test-time adaptation (CTTA) is to adapt a fixed discriminative pre-trained source model, parameterized by ${\theta_{s}}$ and trained on the source dataset ($\mathcal{X}^{source}$, $\mathcal{Y}^{source}$), to multiple incoming tasks $\mathcal{T}_{i}$ = {$x_{k}$}${}_{k=1}^{n}$ where $\mathcal{T}_{i}$ refers to the $i^{th}$ task and n is the total number of batches of $\mathcal{T}_{i}$, with changes in domain distributions between tasks. At each time step, ${\theta_{s}}$ is gradually adapted to each incoming batch ($x_{k}$) of task $\mathcal{T}_{i}$ in an online manner, where the model parameters are updated to ${\theta}$. In this entire process of adaptation, the source dataset is unavailable for use, due to privacy and storage constraints. As evident, the source model is first parameterized by ${\theta_{s}}$ before the adaptation begins. With the continuum of tasks, at time step t, the complete model parameters ${\theta}^{t}$ consists of a matrix of parameters, given by ${\theta}^{t}$ = ${\theta^{t}_{1}}$, $\theta^{t}_{2}$,…, ${\theta^{t}_{n}}$,…., ${\theta^{t}_{N}}$}, where ${\theta^{t}_{n}}$, at time t, denotes the parameters of the $n^{th}$ trainable layer out of a total of N.
3.2 Analysis of Adaptive Learning Rate Schemes for CTTA
In this subsection, we provide a detailed view and analysis of surgical fine-tuning [21], AutoLR [35], and LAW [33]. This is essential to understand the design principles and reasons for having an adaptive learning rate scheme for CTTA, which also includes the motivation of our work.
To preserve the information during the source pre-training phase while fine-tuning on a target task of a different distribution, surgical fine-tuning [21] empirically show that heuristically fine-tuning only the first convolutional block results in a better performance over full fine-tuning. Even though it is a desirable outcome for efficient fine-tuning, in this work, we close the gap by automatically selecting the layers that would adapt well to the target task.
Inspired by [47, 49], AutoLR [35] claims that the weight changes across layers should be in increasing order, where the weight change $\delta w_{n}$ in the $n^{th}$ layer with weights $w_{n}$, at epoch e, is defined as,
$\delta w_{n}=\|w_{n}^{e}\ -w_{n}^{(e-1)}\|$ (1)
They hypothesize that weight changes in the initial layers should be small because they extract general features whereas the deeper layers extract task-specific features. A constrained adaptive LR optimization scheme is set up to force the weight variations to increase with the layers without studying the importance of each layer during adaptation. However, the updated LRs, for each layer, do not deviate too much and remain constant throughout.
As the first adaptive LR work for CTTA, LAW [33] hypothesizes that identifying and controlling the adaptation rate of the different layers is useful [21]. To understand the loss landscape and layer importance, the authors compute the gradient from the log-likelihood as,
$s(\theta_{n}^{t};x_{k})=\nabla_{\theta_{n}^{t}}\log(p_{\theta^{t}}(x_{k}))$ (2)
where $x_{k}$ and $\theta_{n}^{t}$ refer to the current test batch and parameters of layer n at time t. Then, the Fisher Information Matrix (FIM) [36], as a domain-shift indicator, is computed as a first-order approximation of the Hessian matrix. As it can be seen, the layer-wise importance (Eq. 2) completely relies on the pseudo-labels to compute the log-likelihood, which is highly unreliable and uncertain. Due to the distribution shift, the label quality would begin to decrease for a continuum of tasks. Furthermore, the authors propose to capture "domain-level" FIM ($\hat{F_{n}^{t}}$), for each layer during the adaptation, by a direct accumulation of the current layer-wise FIM (${F_{n}^{t}}$) as,
$\hat{F}_{n}^{t}=\hat{F}_{n}^{t-1}+F_{n}^{t}$ (3)
Based on Eq. 3, this means that LAW [33] puts equal emphasis on $\hat{F}_{n}^{t}$ and ${F_{n}^{t}}$, for each layer. An issue arises when ${F}_{n}^{t}$ of the current batch is small but gets outweighed by a "relatively" larger $\hat{F}_{n}^{t}$. This raises questions on the amount of adaptation required by the $n^{th}$ layer for batch $x_{k}$ as the FIMs dynamically change for each test batch. This leads to an imbalance between the accumulated FIMs $\hat{F}_{n}^{t}$ and the current FIM ${F_{n}^{t}}$.
Drawing from these primary observations, our method focuses on addressing the following: 1) Based on the prediction uncertainty of the model, we automatically select a subset of layers that would require further adaptation to a task, without relying on the pseudo-labels. 2) As an indicator of the domain shift, we leverage the idea of parameter sensitivity of these layers only and adaptively alter their learning rates. We freeze the layers that are not selected.
4 Proposed Method
Figure 1: PALM framework: At time t, input batch $x_{k}$ is processed by the model, parameterized by $\theta^{t}$. The KL-divergence between the softmax predictions and a uniform distribution is backpropagated to select the layers with the gradient norm $\leq$ $\eta$, to quantify the uncertainty. The parameter sensitivities of these layers, as an indicator of domain shift meters, are computed to update their learning rates (Eq. 11). Finally, with the optimization objective in Eq. 16, we update the model with the adjusted learning rates of the parameters.
In this section, we introduce and formalize our proposed method for CTTA to address the key limitations of the works discussed above. Fig. 1 diagrammatically illustrates our entire approach.
4.1 Layer Selection by Quantifying Prediction Uncertainty
CTTA maximizes model performance in the current domain and is vital to assess prediction certainty amidst domain shifts. In the literature on continual learning and TTA, several works utilize FIM [18, 38, 33]. However, the unavailability of true labels, with an unknown distribution, motivates us to capture the prediction uncertainty. Prior literature proposed methods for uncertainty from the classifier output or feature space [12, 14, 24, 25, 20]. Motivated by [15], which proposed an uncertainty scoring function from the gradient space, we compute the gradients of each layer. In essence, we backpropagate the KL divergence (KL) between the softmax probabilities and a uniform distribution u = [$\frac{1}{C}$, $\frac{1}{C}$, …., $\frac{1}{C}$] $\in$ $\mathbb{R}^{C}$, where C is the total number of classes. This is a measurement of how far the distribution of the model’s prediction is from a uniform distribution. Intuitively, due to the distribution shift, the KL divergence should be smaller. Let h denote the classifier of the model parameterized by ${\theta^{t}}$. We compute the loss, denoting it as $\mathcal{L}(\theta^{t})$ at time stamp t, as follows,
$\mathcal{L}(\theta^{t})=\text{KL}(\text{softmax}(\hat{h}(x_{k}))||\text{{u}})$ (4)
where $\hat{h}(x_{k})$ is the classifier’s output logits ($h(x_{k})$) smoothed by a temperature coefficient T i.e. $\hat{h}(x_{k})$ = $h(x_{k})$/T. Our intuition is that batches within the same task are of the same distribution and so, smoothing the logits with large values of T will lead to them being indistinguishable. With model parameters $\theta^{t}$, the gradients (${{\nabla\mathcal{L}({\theta}^{t})}}$ = $\frac{\partial\mathcal{L}(\theta^{t})}{\partial{\theta}^{t}}$) can be simplified to,
${{\nabla\mathcal{L}({\theta}^{t})}}=\frac{1}{C}\sum_{j=1}^{C}\frac{\partial\mathcal{L_{\text{CE}}}(\hat{h}(x_{k}),j)}{\partial\theta^{t}}$ (5)
On the whole, the gradients boil down to the averaged cross-entropy loss between the smoothed output logits and a uniform label (Eq. 5). In test-time adaptation, since ground-truth labels are not present, the gradients are more indicative of familiarity of the model with the batch of data. For batches within the same task, that share the same distribution, the vector norm of Eq. 5 would be higher because of a higher KL divergence since the softmax predictions would be less uniformly distributed. However, if the norm is smaller, this would mean that the softmax outputs are closer to u. Our key idea is to select all the parameters of layers that have the norm of gradients to be less than or equal to a threshold $\eta$. Layers with small gradient norms would require more adaptation for the model to make a confident prediction. Since we freeze the remaining layers, this also helps to alleviate catastrophic forgetting to a large extent by preserving the prior knowledge. And so the scoring function, for each layer, is defined as,
$\mathcal{Z}_{\theta_{n}}=\|{{\nabla\mathcal{L}({\theta}_{n}^{t})}}\|_{1}$ (6)
As mentioned earlier, $\mathcal{Z}_{\theta_{n}}$leq $\eta$ allows the model to adaptively determine which specific layers need further adaptation. We freeze the layers with $\mathcal{Z}_{\theta_{n}}$ $>$ $\eta$ i.e. set their learning rate to 0. This way, in contrast to surgical fine-tuning [21], we automatically select a subset of layers for adaptation to a batch of data.
4.2 Parameter Sensitivity as an Indicator of Domain Shift
New studies have found that model parameters demonstrate levels of criticality [2, 50] across different positions. Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors [31, 34]. Inspired by this observation, we claim that parameters have different contributions during test-time adaptation. From Sec. 4.1, we adaptively select layers that are uncertain in their predictions. For parameters of these layers only, we compute their sensitivity i.e. an approximation of an error when it is removed.
In our work, motivated from [19, 30, 29], we compute the parameter sensitivity of a parameter $\theta_{n}^{t}$ based on the loss (Eq. 4) as follows,
$S_{n}^{t}=\mathcal{L}({\theta}^{t})-\mathcal{L}({\theta}^{t}|\theta_{n}^{t}=0)$ (7)
Precisely, $S_{n}^{t}$ approximates the change in the loss landscape at $\theta^{t}$ when $\theta_{n}^{t}$ is removed. Sensitivity, computed as a "weighting factor" of model weights [30, 29, 23], determines the level of adaptation needed for model parameters.
We utilize the gradients of parameters from the selected layers, computed by Eq. 5 to compactly derive Eq. 7 from the first-order Taylor series expansion of $\mathcal{L}(\theta^{t})$ as,
$S_{n}^{t}=|\theta_{n}^{t}\nabla\mathcal{L}({\theta_{n}}^{t})|$ (8)
Our rationale is as follows: parameters with low sensitivity are pushed to have higher learning rates owing to more adaptation. However, $S_{n}^{t}$ of $\theta_{n}^{t}$ can be highly uncertain and would have a large variation. We introduce a method to capture this "domain-level" uncertainty by computing a weighted moving average of the sensitivity $S_{n}^{t}$. This is formulated as,
$\hat{S}_{n}^{t}=\alpha S_{n}^{t}+(1-\alpha)\hat{S}_{n}^{(t-1)}$ (9)
where $\hat{S}_{n}^{t}$ denotes the "domain-level" sensitivity and $\alpha$in [0,1] is a smoothing factor. Specifically, $\alpha$ decides the contribution of the current sensitivity $S_{n}^{t}$ to $\hat{S}_{n}^{t}$. Since CTTA happens at batch level, it is critical to achieve a good balance between $S_{n}^{t}$ and $\hat{S}_{n}^{t}$, with a slightly larger focus on $S_{n}^{t}$. As it can be seen, we do not directly accumulate the "domain-level" information, unlike LAW [33].
In addition, due to large distribution shifts happening because of the continuum of tasks, larger deviations of $S_{n}^{t}$ from $\hat{S}_{n}^{t}$ are bound to happen. To quantify these statistics, an absolute difference between $\hat{S}_{n}^{t}$ and $S_{n}^{t}$ can be computed as,
$\hat{D}_{n}^{t}=|S_{n}^{t}-\hat{S}_{n}^{t}|$ (10)
We emphasize that since Eq. 9 leverages sensitivity from previously computed batches and domains, $\hat{D}_{n}^{t}$ would act as a good indicator of how uncertain $S_{n}^{t}$ is to the previous batches and domains, for $\theta_{n}$. This would take care of error accumulation too.
To automatically alter the learning rates of these parameters, the learning rate "importance" ($i_{n}^{t}$) is calculated as follows,
$i_{n}^{t}=(\hat{D}_{n}^{t})/(\hat{S}_{n}^{t})$ (11)
Then, the final learning rate to update $\theta_{n}^{t}$ would be
$\hat{\kappa}_{n}^{t}=\kappa*i_{n}^{t}$ (12)
where $\kappa$ is the base fixed learning rate. The intuition behind the formulation of Eq. 11 is as follows - smaller values of $\hat{S}_{n}^{t}$ indicate lower sensitivity and hence larger parameter updates are required via promoting the learning rate. At the same time, large uncertainties by $\hat{D}_{n}^{t}$ would also require larger updates. To avoid any overflow/underflow, we add $\epsilon$ (0<$\epsilon$$\ll$1) to the numerator and the denominator in Eq. 11.
4.3 Optimization Objective
Though our method is robust, it is necessary to avoid miscalibrated labels. TENT [44] minimized the Shannon entropy on the output logits, defined as,
$\mathcal{H}(\hat{y})=-\sum_{c}p(y_{c})log(p(y_{c}))$ (13)
where $p(y_{c})$ refers to the softmax probability of class c i.e. $p(y_{c})$ = softmax($h(x_{k})$). Following previous CTTA work [41], we use entropy minimization as the main loss objective for adaptation where Eq. 13 is minimized for samples below a certain threshold.
$\mathcal{L}^{entropy}=\mathds{1}_{\mathcal{H}(\hat{y})\leq\mathcal{H}_{0}}\mathcal{H}(\hat{y})$ (14)
where $\mathds{1}$ is the indicator function and $\mathcal{H}_{0}$ is a fixed threshold set to $0.4*lnC$, consistently following [41]. To be fair with prior works [45, 33], we use the consistency loss between the model predictions on $x_{k}$ and its augmented version $\hat{x}_{k}$ to regularize the model as follows,
$\mathcal{L}^{const}=-\sum_{c}p(y_{c})log(p(\hat{y}_{c}))$ (15)
where, $\hat{y}_{c}$ = $h(\hat{x}_{k})$. Our overall optimization objective is as follows,
$\mathcal{L}^{total}=\mathcal{L}^{entropy}+\lambda\mathcal{L}^{const}$ (16)
where $\lambda$ indicates a regularisation coefficient to balance the consistency loss. In summary, upon calculating the parameter-specific learning rates (Eq. 12), we optimize the continually fine-tuned model to minimize Eq. 16.
5 Experiments
We draw a contrast of our proposed method against other state-of-the-art works in a continual test-time adaptation (CTTA) setting. However, to have a broader study and analysis, we also report results in a gradual test-time adaptation (GTTA) setting, following [45, 33]. In the following subsections, we describe the implementation details of both the challenging adaptation settings. For additional results and analysis, please refer to the Supplementary Material.
5.1 Experimental Settings and Baselines
Datasets and tasks. Following the standard benchmarks in CTTA, set by CoTTA [45], we evaluate our proposed method on datasets - CIFAR-10C, CIFAR-100C, and ImageNet-C, based on image corruption schemes as set in [11]. Each dataset contains a set of 15 corruption styles (e.g. gaussian noise, shot noise, impulse noise, defocus blur) with 5 severity levels. In line with established practices outlined in [45, 6], we report the mean classification error as the evaluation metric.
CTTA implementation details. We evaluate our approach using frameworks from prior state-of-the-art methods [44, 45, 3, 41, 37, 33]. To maintain fairness regarding the usage of pre-trained vision models, we employ WideResNet-28 [48] for CIFAR-10C, ResNeXt-29 [46] for CIFAR-100C, and ResNet-50 [10] for ImageNet-C, all available on RobustBench [5]. We optimize using an Adam optimizer [17], setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C. For balancing parameter sensitivities, we set $\alpha$ to 0.5, 0.9, and 0.5 respectively, with temperature coefficients T set to 50, 100, and 1000 respectively. We set $\eta$ to 1, 0.5, and 0.3, and $\lambda$ to 0.01 throughout. Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA [45], and results are reported for the highest severity level of 5 for each task.
GTTA implementation details. In the standard CTTA setup, each task receives images with the highest severity. However, in GTTA, following [45, 41, 33], for each task the severity levels are gradually changed as …3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5$\rightarrow$4$\rightarrow$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3… We maintain consistency with the experimental details outlined in the CTTA setting.
Baselines. We compare our work against several CTTA baselines, including Source, BN Stats Adapt [37], TENT Continual [44], CoTTA [45], EcoTTA [41], Adacontrast [3], and LAW [33]. In addition, we also contrast our work against surgical fine-tuning [21]. For EcoTTA, we report the results on CIFAR-10C and ImageNet-C only since they do not use ResNeXt-29 [46] for CIFAR-100C. Following the findings of surgical fine-tuning [21], we continually fine-tune only the first convolutional block of the respective models, denoted as Surgical fine-tuning (continual). We have mentioned the training details in the Supplementary. For the reported results of LAW [33], we reproduce their results strictly following their official implementation  [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
Table 1: Mean errors (in %) on CIFAR-10C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5. We report the error of a single update (online) only using WideResNet-28 [48] for all the reported baselines.
Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
Elastic
Transform
Pixelate JPEG Mean
Source 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.5 30.3 43.5
BN Stats Adapt [37] 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
TENT Continual [44] 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
CoTTA [45]
24.3 21.3 26.6 11.6 27.6 12.2 10.3 14.8 14.1 12.4 7.5 10.6 18.3 13.4 17.3 16.2
ECoTTA [41]
23.8 18.7 25.7 11.5 29.8 13.3 11.3 15.3 15.0 13.0 7.9 11.3 20.2 15.1 20.5 16.8
Adacontrast [3]
29.1 22.5 30.0 14.0 32.7 14.1 12.0 16.6 14.9 14.4 8.1 10.0 21.9 17.7 20.0 18.5
Surgical
fine-tuning [21]
27.71 24.14 33.79 12.18 33.07 13.09 10.97 15.94 15.61 13.8 7.84 10.77 21.17 16.62 22.88 18.64
LAW [33]
24.71 18.91 25.56 12.96 26.78 15.09 11.8 15.17 14.78 15.89 10.13 13.89 19.41 14.74 18.35 17.21
Ours
25.89 18.1 22.71 12.38 25.3 13.18 10.79 13.5 13.15 12.21 8.53 11.85 17.98 12.02 15.47 15.54
Table 2: Mean errors (in %) on CIFAR-100C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5. We report the error of a single update (online) only using ResNeXt-29 [46] for all the reported baselines.
Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
Elastic
Transform
Pixelate JPEG Mean
Source
73.0 68 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
BN Stats Adapt [37]
42.1 40.7 42.7 27.6 41.9 29.7 27.9 34.9 35.0 41.5 26.5 30.3 35.7 32.9 41.2 35.4
TENT Continual [44]
37.2 35.8 41.7 37.9 51.2 48.3 48.5 58.4 63.7 71.1 70.4 82.3 88.0 88.5 90.4 60.9
CoTTA [45]
40.1 37.7 39.7 26.9 38.0 27.9 26.4 32.8 31.8 40.3 24.7 26.9 32.5 28.3 33.5 32.5
Adacontrast [3]
42.3 36.8 38.6 27.7 40.1 29.1 27.5 32.9 30.7 38.2 25.9 28.3 33.9 33.3 36.2 33.4
Surgical
fine-tuning [21]
40.75 36.36 37.89 25.53 39.02 27.65 25.27 32.26 30.45 37.35 23.78 26.82 32.37 27.64 37.20 32.02
LAW [33]
41.06 36.73 38.38 25.67 37.02 27.82 25.22 30.79 30.04 37.35 24.41 27.69 31.25 27.84 34.92 31.75
Ours
37.33 32.5 34.94 26.20 35.28 27.54 24.64 28.85 29.15 34.14 23.52 26.99 31.15 26.56 34.1 30.19
5.2 CTTA Classification Results
Results on CIFAR-10C and CIFAR-100C. We contrast our proposed method against several baselines including surgical fine-tuning [21]. Tables 1 and 2 report the CTTA mean classification errors across all the domains, for each of the methods. The results reveal that our method performs the best across all training setups and baselines for both datasets. In particular, we achieve large improvements of 4.86% and 5.16% for CIFAR-10C and 5.21% and 30.71% for CIFAR-100C, over BN Stats Adapt[3] and TENT Continual[44] respectively. Concerning recent work LAW [33] that uses an adaptive LR scheme, our method achieves an improvement of 1.67% and 1.56% on CIFAR-10C and CIFAR-100C, respectively. It’s noteworthy that TENT Continual initially excels in the early tasks but experiences error accumulation over time, leading to an increase in mean error. In the study by CoTTA [45], despite achieving good performance, the approach’s reliance on full model updates renders it computationally expensive and unsuitable for deployment. LAW [33] uses the uncertain pseudo labels, of each batch, to calculate the domain shift approximated via the FIM computation. The minor error that is introduced for each batch begins to propagate across all the domains. In addition, the imbalance caused by the "domain-level" FIM and the current batch’s FIM, results in a higher error rate for the initial tasks. In comparison, the computation of the model prediction uncertainty to select layers is useful in our work (Eq. 4). Another interesting observation is that with a drastic reduction in the number of trainable parameters, as in Surgical fine-tuning (continual), the mean errors in Table 2 is better than most CTTA benchmarks, if not the best. This affirms the findings in [21] that concentrated adaptation, for image corruptions, in the initial layers is sufficient.
Results on Imagenet-C In Table 3, we report the CTTA results on the ImageNet-C dataset, with ResNet-50 pre-trained on ImageNet. As shown, our method outperforms all the previous CTTA baselines and Surgical fine-tuning (continual). We observe an improvement of 1% over LAW [33]. In addition, we achieve 2.31%, 5.38%, and 30.71% improvements over CoTTA [45], Adacontrast [3], and TENT [44], respectively.
Table 3: Mean errors (in %) on ImageNet-C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5. We report the error of a single update (online) only using ResNet-50 [10] for all the reported baselines.
Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
Elastic
Transform
Pixelate JPEG Mean
Source
97.8 97.1 98.2 81.7 89.8 85.2 78.0 83.5 77.1 75.9 41.3 94.5 82.5 79.3 68.6 82.0
BN Stats Adapt [37]
85.0 83.7 85.0 84.7 84.3 73.7 61.2 66.0 68.2 52.1 34.9 82.7 55.9 51.3 59.8 68.6
TENT Continual [44]
81.6 74.6 72.7 77.6 73.8 65.5 55.3 61.6 63.0 51.7 38.2 72.1 50.8 47.4 53.3 62.6
CoTTA [45]
84.7 82.1 80.6 81.3 79.0 68.6 57.5 60.3 60.5 48.3 36.6 66.1 47.2 41.2 46.0 62.7
ECoTTA [41]
- - - - - - - - - - - - - - - 63.4
Adacontrast [3]
82.9 80.9 78.4 81.4 78.7 72.9 64.0 63.5 64.5 53.5 38.4 66.7 54.6 49.4 53.0 65.5
Surgical
fine-tuning [21]
83.42 79.5 77.28 86.1 82.4 72.28 61.18 63.1 64.76 51.26 35.92 72.9 56.24 46.72 50.2 65.55
LAW [33]
81.86 75.04 72.12 77.42 73.22 63.96 54.6 57.92 61.06 50.1 36.36 68.66 49.08 46.22 49.12 61.12
Ours
81.06 73.32 70.88 77.00 71.88 62.34 53.94 56.7 60.76 50.36 36.26 65.96 48.08 45.28 48.0 60.12
Table 4: Mean errors (in %) on CIFAR-10C, CIFAR-100C, and ImageNet-C - We report the gradual TTA (GTTA) mean errors averaged across all the severity levels (1normal-$\rightarrow$2normal-$\rightarrow$3normal-$\rightarrow$4normal-$\rightarrow$5).
Dataset Source
BN Stats
Adapt
TENT
Continual
CoTTA Adacontrast
Surgical
fine-tuning
LAW Ours
CIFAR-10C 24.7 13.7 20.4 10.9 12.1 11.33 10.87 9.56
CIFAR-100C 33.6 29.9 74.8 26.3 33.0 25.31 27.19 26.11
ImageNet-C 58.4 48.3 46.4 38.8 66.3 41.36 40.52 38.49
5.3 GTTA Classification Results
Results on CIFAR-10C, CIFAR-100C, and Imagenet-C. This subsection presents the gradual test-time adaptation (GTTA) results on CIFAR-10C, CIFAR-100C, and Imagenet-C. For each task, the corruption severity of the inputs is gradually increased from levels 1 to 5, as discussed in the implementation details in Sec. 5.1. Table 4 reports the mean classification error for GTTA across all severities. We observe that, for all the datasets, our approach outperforms or performs comparably well across all the baselines [37, 44, 45, 3, 33]. With most parameters frozen during the adaptation on each batch, our method is more robust and inexpensive than CoTTA [45], which also demonstrates competitive improvements in the mean classification error.
5.4 Ablation Studies
We conduct ablation experiments to validate our approach and gain insights into its underlying mechanisms and components. Specifically, we conduct these experiments and report results on the three CTTA benchmark datasets: CIFAR-10C, CIFAR-100C, and ImageNet-100C.
    WideResNet-28 [48]       ResNeXt-29 [46]       ResNet-50 [10]
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Illustrations of the learning rate importances of WideResNet-28 [48], ResNeXt-29 [46], and ResNet-50 [10] during CTTA, for different convolutional blocks/stages/layers. [Top row] - Variations across respective datasets for domain glass_blur. [Bottom row] - Variations across respective datasets for domain snow.
Analysis of the learning rate importance. The prime objective of our work is to adaptively alter the learning rates of parameters of selected layers based on their sensitivity. In Fig. 2, we illustrate how the learning rate (LR) importance (Eq. 11) of WideResNet-28 [48], ResNeXt-29 [46], and ResNet-50 [10], averaged for a task, change during adaptation on their respective datasets. For brevity, we show two domains - glass_blur and snow. As we can see, the importance varies across domains and models. Across all the models, we notice that generally, the deeper layers have smaller associated LR importances, which means that the initial layers are adapted more to the target domain. This result completely aligns with the findings of surgical fine-tuning [21]. In addition to that, we also notice more sparsity in the deeper layers. This further confirms the necessity of freezing parameters automatically selected based on the target domain for adaptation.
Nullifying prediction uncertainty. In LAW [33], the authors make use of the pseudo-labels to automatically decide on the importance of layers for adaptation. In this, we study the effectiveness of such a design for our method i.e. we investigate if layer importance computed using pseudo-labels is helpful. We compute the gradients from the log-likelihood (Eq. 2) and estimate the sensitivity of the parameters of the chosen layers. Table 5.4 reports the mean classification errors for all the datasets. We observe that the gradients computed using the pseudo-labels are an unreliable indicator of prediction uncertainty. The scoring function (Eq. 6) computed using these gradients carries insufficient information. We achieve improvements of 2.44%, 3.69%, and 8.5% over CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
Refer to caption Refer to caption
a) Ablation on smoothing factor $\alpha$. b) Ablation on temperature T.
Figure 3: Ablation results on smoothing factor $\alpha$ and temperature T for CIFAR-10C, CIFAR-100C, and ImageNet-C.
Table 5: Prediction uncertainty using pseudo-labels vs our method. Dataset Pseudo-labels Ours CIFAR-10C 17.98 15.54 CIFAR-100C 33.88 30.19 ImageNet-C 68.62 60.12
Table 6: Ablation results on benchmark datasets with varying $\lambda$. $\lambda$ 0 0.01 0.05 0.1 0.5 1.0 CIFAR-10C 16.84 15.54 16.45 17.03 17.65 17.93 CIFAR-100C 30.44 30.19 30.51 30.57 30.63 30.7 ImageNet-C 63.98 60.12 60.38 60.4 60.4 60.42
Contribution of smoothing factor $\alpha$. An important component of our method is the computation of the moving average of sensitivities across data batches, smoothed by a parameter $\alpha$ (Eq. 9). Fig. 3(a) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$. We observe that the "domain-level" sensitivity across past batches of data is indeed useful. For $\alpha$ = 0 i.e. when the current sensitivity of parameters of selected layers is not considered, the mean classification error for the respective datasets drastically spikes up. This suggests that the effect of the current sensitivity of selected layers is indeed required. For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.67% over LAW [33]. Similarly, for CIFAR-100C and ImageNet-C, we see improvements of 1.56% and 1% over LAW [33], at $\alpha$ = 0.9 and 0.5 respectively. With higher values of $\alpha$, the results do not vary too much and are stable. Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{n}^{t}$) and current sensitivity ($S_{n}^{t}$) is critical, as seen in the figure.
[…]
Robustness to the regularisation coefficient $\lambda$. As an important ablation step, we understand the effect of the parameter $\lambda$ on the final optimization objective, as in Eq. 16. We set $\lambda$in {0, 0.01, 0.05, 0.1, 0.5, 1.0}. We study the contributions of the consistency loss to entropy minimization. Table 5.4 reports the average classification errors on CIFAR-10C, CIFAR-100C, and ImageNet-C, with varying $\lambda$. For $\lambda$ = 0.01, we achieve the best results across all the datasets. In addition, we observe that our method is very robust to $\lambda$ i.e. with subtle variations to differing $\lambda$, across the datasets. In particular, for CIFAR-10C, with $\lambda$ = 0 i.e. with no consistency loss, we achieve improvements of 0.37% and 3.86% over LAW [33] and TENT Continual [44]. For CIFAR-100C, with no consistency loss ($\lambda$ = 0), we notice improvements of 1.31%, 2.06%, 4.96%, and 30.46% over LAW [33], CoTTA [45], BN Stats Adapt [37] and TENT Continual [44]. As for ImageNet-C, the consistent best setting of $\lambda$ = 0.01 is better than LAW [33] by 1% and by 5.38% over Adacontrast [3]. However, with $\lambda$ = 0 i.e. no consistency loss, we see a spike in the error rate by about 3.86% over the best setting ($\lambda$ = 0.01).
6 Conclusion
In this paper, we propose a novel approach called PALM, to continual test-time adaptation. Our approach automatically decides on the importance of a layer by leveraging information from the gradient space i.e. by backpropagating the KL divergence between a uniform distribution and the softmax outputs. This effectively eliminates the need for pseudo-labels. For the parameters of these layers, selected based on the norm of the gradients, the learning rates are adjusted based on their sensitivity to domain shift. Extensive experiments are conducted on benchmark datasets to demonstrate the efficacy of our method against different baselines. The results reveal that PALM achieves superior performance and pushes the state-of-the-art on several standard baselines.

### v2

@@
 Real-world vision models in dynamic environments face rapid shifts in domain distributions, leading to decreased recognition performance.
-Continual test-time adaptation (CTTA) directly adjusts a pre-trained source discriminative model to these changing domains using test data.
-A highly effective CTTA method involves applying layer-wise adaptive learning rates, and selectively adapting pre-trained layers.
+Using unlabeled test data, continual test-time adaptation (CTTA) directly adjusts a pre-trained source discriminative model to these changing domains.
+A highly effective CTTA method involves applying layer-wise adaptive learning rates for selectively adapting pre-trained layers.
 However, it suffers from the poor estimation of domain shift and the inaccuracies arising from the pseudo-labels.
-In this work, we aim to overcome these limitations by identifying layers through the quantification of model prediction uncertainty without relying on pseudo-labels.
+This work aims to overcome these limitations by identifying layers for adaptation via quantifying model prediction uncertainty without relying on pseudo-labels.
 We utilize the magnitude of gradients as a metric, calculated by backpropagating the KL divergence between the softmax output and a uniform distribution, to select layers for further adaptation.
-Subsequently, for the parameters exclusively belonging to these selected layers, with the remaining ones frozen, we evaluate their sensitivity in order to approximate the domain shift, followed by adjusting their learning rates accordingly.
-Overall, this approach leads to a more robust and stable optimization than prior approaches.
-We conduct extensive image classification experiments on CIFAR-10C, CIFAR-100C, and ImageNet-C and demonstrate the efficacy of our method against standard benchmarks and prior methods.
+Subsequently, for the parameters exclusively belonging to these selected layers, with the remaining ones frozen, we evaluate their sensitivity to approximate the domain shift and adjust their learning rates accordingly.
+We conduct extensive image classification experiments on CIFAR-10C, CIFAR-100C, and ImageNet-C, demonstrating the superior efficacy of our method compared to prior approaches.
 1 Introduction
@@
 While this approach effectively adapts the model to a single type of corruption, it faces two constraints when continually applied across a sequence of test tasks featuring different corruptions.
-Firstly, there is the issue of error accumulation: since pseudo labels are generated for unlabeled test data, errors accumulate over time.
-Secondly, catastrophic forgetting occurs, resulting in a long-term loss of the model’s pre-trained knowledge due to continuous adaptations and large parameter updates across tasks.
+Firstly, there is the issue of error accumulation since pseudo labels are generated for unlabeled test data, errors accumulate over time.
+Secondly, catastrophic forgetting occurs, resulting in a long-term loss of the model’s source knowledge due to adaptations and large parameter updates across tasks.
 Continual test-time adaptation (CTTA) has been recently introduced to tackle these challenges.
@@
 CoTTA [37] includes full model updates, with a teacher-student network setup, to handle the above issues.
-To further improve the performance of CoTTA, EcoTTA [41] introduces the attachment of "meta networks", i.e., a BN layer and a convolution block, to the frozen source model for efficient adaptation.
-More recently, the authors in [21] demonstrated that based on the type of distributional shift of the test data, a certain subset of layers of a pre-trained model can be "surgically" or heuristically chosen to fine-tune.
+To further improve the performance of CoTTA, EcoTTA [33] introduces the attachment of “meta networks", i.e., a BN layer and a convolution block, to the frozen source model for efficient adaptation.
+More recently, the authors in [15] demonstrated that based on the type of distributional shift of the test data, a certain subset of layers of a pre-trained model can be “surgically" or heuristically chosen to fine-tune.
 Inspired by this finding, Layer-wise Auto-Weighting (LAW) [24] was introduced recently.
-LAW utilizes the Fisher Information Matrix (FIM) [36] to estimate domain shifts, automatically weighting trainable parameters and adjusting their learning rates accordingly.
+LAW utilizes the Fisher Information Matrix (FIM) [27] to estimate domain shifts, automatically weighting trainable parameters and adjusting their learning rates (LRs) accordingly.
 While LAW achieves state-of-the-art results for CTTA, it still encounters two significant limitations: 1) It presupposes that layer-wise importance can be accurately estimated via FIM with initially generated pseudo-labels.
@@
 2) Relying on directly accumulating such approximated importance from noisy labels to calculate domain shifts can lead to inaccuracies.
-In this paper, we present a novel method called PALM to push the limit of adaptive learning rate methods for CTTA.
-Our approach begins by identifying a subset of layers that significantly contribute to the uncertainty of model predictions.
-Rather than relying on pseudo-labels, we determine layer-wise importance by leveraging information from the gradient space, specifically by calculating the KL divergence between the model output and a uniform distribution.
+In this paper, we present a novel method called PALM to push the limits of adaptive learning rate methods for CTTA.
+To address catastrophic forgetting, our approach begins by identifying a subset of layers that significantly contribute to the uncertainty of model predictions.
+Rather than relying on pseudo-labels, we determine layer-wise importance by leveraging information from the gradient space, specifically by calculating the Kullback-Leibler (KL) divergence between the model output and a uniform distribution.
 This approach eliminates dependence on pseudo-labels and enables a more accurate estimation of layer-wise importance.
-Subsequently, the learning rates of the parameters of these layers are adjusted based on the degree of domain shift observed in the test data.
-We utilize parameter sensitivity as an indicator of the degree of distribution shift and employ weighted moving averages to aggregate sensitivity from previous data batches.
-Compared to existing adaptive learning rate methods, our proposed approach offers a better estimation of domain shift for adjusting learning rates.
-To summarise, our contributions are as follows:
-  • We present PALM, a novel approach to continual test-time adaptation. Specifically, we demonstrate improved schemes for adaptive learning rate methods for CTTA.
-  • Considering the prediction uncertainty of the model at test-time, we design an approach to automatically select layers that would require further adaptation and push the learning rates of parameters of these layers for model updates, by estimating the domain shift based on parameter sensitivity.
-  • Our method delivers promising results in both continual and gradual TTA across three standard benchmark datasets.
-In addition, we conduct and demonstrate ablation experiments on the different components of our work.
+Subsequently, the LRs of the parameters of these selected layers are adjusted based on the degree of domain shift observed in the test data.
+We use parameter sensitivity as an indicator of domain shift and employ weighted moving averages to aggregate sensitivity from previous test batches.
+Our approach provides a more accurate estimation of domain shift compared to existing adaptive LR methods, leading to more effective adjustments in LRs.
+To summarise, our contributions are as follows: 1) We present PALM, introducing novel improved schemes for CTTA with a specific focus on addressing the key limitations of existing adaptive learning rate methods.
+2) Considering the model’s prediction uncertainty at test-time, we automatically select and adapt layers requiring further updates while freezing the rest.
+We adjust learning rates based on parameter sensitivity to estimate domain shift, making the overall process computationally inexpensive.
+3) Our method delivers superior results, via extensive evaluations, in both continual and gradual TTA across three standard benchmark datasets.
 2 Related Works
-Test-time adaptation (TTA): The overarching target of TTA is to adapt source models to unlabelled test data in real-time in a source-free online training fashion [16, 42, 44].
-TENT [44] optimizes the Batch Normalisation (BN) parameters by minimizing the Shannon entropy [40] of the pseudo labels.
-In [37], the statistics of the BN parameters are replaced by the ones of the input corrupted data.
-Adacontrast [3] combines contrastive learning and utilizes the memory module to update the entire model’s parameters.
-While previous TTA methods showed significant improvements, they were demonstrated for a single domain shift at a time [4, 26, 37, 44] and did not consider continual distribution shifts.
-Continual TTA (CTTA): CTTA poses an adaptation challenge wherein the model must continually adjust to evolving target data, accommodating a sequence of dynamic domain shifts.
-The first work, CoTTA [45], considers addressing this challenging task by setting up a teacher-student network, with full model updates.
-The teacher model is updated using an exponential moving average (EMA) to deal with forgetting and accumulation of errors with time.
-EcoTTA [41] introduces "meta networks" i.e.
-a BN layer and a convolution block are attached to the frozen source model.
-At every stage, an $L_{1}$ regularisation loss between the outputs of the meta network and that of the source model is computed to handle catastrophic forgetting.
-Efficient fine-tuning approaches: Freezing model parameters to learn information from downstream data has witnessed a good body of work [39, 27, 51, 22, 7, 9].
-Depending upon the distribution shift of the target task, in [21], the authors demonstrated that heuristically selecting a subset of convolutional blocks for training while freezing the remaining blocks works better than full fine-tuning.
-Adaptive-learning rate methods: Regarding the studies on the importance of each layer in training, [47, 49] concluded that deeper layers extract task-specific features, as opposed to the shallow layers.
-Based on this finding, AutoLR [35] design an adapted learning rate (LR) scheme based on the weight variations between layers.
-Despite attempts to set different initial learning rates for each layer, these rates remain mostly constant throughout training.
-In the context of CTTA, LAW [33] used the FIM [36] to calculate the domain shift and automatically adjust the learning rates of parameters.
-However, to obtain the layer importance, the pseudo-labels are used to compute the FIM, which would be inaccurate.
-In addition, a direct accumulation of FIMs across domains would lead to error accumulation.
-Furthermore, the LR "weight" of each layer is upper-bounded by the base LR that is used, which ignores the relative "sensitivity" of parameters towards adaptation of the current batch and domain.
-Parameter sensitivity: In the literature about neural network pruning, understanding the sensitivity of parameters towards training has been studied well [19, 28, 43, 29, 23].
-The overall idea is that parameters with lower sensitivity need to be adapted more with an increased learning rate, and vice-versa.
-[23], in addition, introduces a temporal variation of sensitivity, due to uncertainty, as another factor during the training of models for natural language understanding and machine translation tasks.
-3 Background
+Test-time adaptation (TTA): The goal of TTA is to adapt source models to unlabelled test data in real-time using a source-free online approach [12, 34, 36].
+TENT [36] optimizes BN parameters by minimizing the Shannon entropy [31] of pseudo labels.
+BN Stats Adapt (BN-1) [28] replaces the BN statistics with those from corrupted test data, while AdaContrast [3] uses contrastive learning and a memory module to update the entire model.
+However, these methods address only single domain shifts and overlook continual shifts.
+Continual TTA (CTTA): CTTA requires models to adapt to dynamic domain shifts.
+CoTTA [37] uses a teacher-student setup with full model updates.
+The teacher model is updated using an exponential moving average (EMA) to deal with catastrophic forgetting and error accumulation with time.
+EcoTTA [33] adds “meta networks" to a frozen source model, computing an $L_{1}$ regularization loss to prevent catastrophic forgetting.
+DSS [38] introduces a dynamic thresholding framework to differentiate between low and high sample quality.
+Efficient model fine-tuning: Freezing model parameters to learn information from downstream data has witnessed a good body of work [30, 17, 45, 7].
+Depending upon the distribution shift of the target task, in [15], the authors demonstrate that heuristically selecting a subset of convolutional blocks for training while freezing the remaining blocks works better than full fine-tuning.
+Adaptive LR methods: Research such as [40, 43] demonstrates that deeper layers of a network extract task-specific features, unlike shallow layers.
+AutoLR [26] developed an adaptive LR scheme based on weight variations between layers, although these rates tend to mostly stay constant during training.
+For CTTA, LAW [24] uses the FIM [27] to assess domain shifts and adjust parameter LRs accordingly.
+However, using pseudo-labels to calculate FIM may lead to inaccuracies, and accumulating FIMs across domains can result in errors.
+Moreover, the adjustment of LR weights for each layer is constrained by the base LR, disregarding the relative “sensitivity" of parameters to the adaptation needs of the current batch and domain.
+Parameter sensitivity: Neural network pruning literature extensively explores parameter sensitivity [14, 18, 35, 20].
+It suggests that less sensitive parameters require higher LRs for effective adaptation, and vice-versa.
+Additionally, [16] introduces a concept of temporal sensitivity variation, influenced by uncertainty, as a factor in training models for tasks like natural language understanding and machine translation.
+3 Preliminary
 3.1 Problem Setting
-The objective of continual test-time adaptation (CTTA) is to adapt a fixed discriminative pre-trained source model, parameterized by ${\theta_{s}}$ and trained on the source dataset ($\mathcal{X}^{source}$, $\mathcal{Y}^{source}$), to multiple incoming tasks $\mathcal{T}_{i}$ = {$x_{k}$}${}_{k=1}^{n}$ where $\mathcal{T}_{i}$ refers to the $i^{th}$ task and n is the total number of batches of $\mathcal{T}_{i}$, with changes in domain distributions between tasks.
-At each time step, ${\theta_{s}}$ is gradually adapted to each incoming batch ($x_{k}$) of task $\mathcal{T}_{i}$ in an online manner, where the model parameters are updated to ${\theta}$.
-In this entire process of adaptation, the source dataset is unavailable for use, due to privacy and storage constraints.
-As evident, the source model is first parameterized by ${\theta_{s}}$ before the adaptation begins.
-With the continuum of tasks, at time step t, the complete model parameters ${\theta}^{t}$ consists of a matrix of parameters, given by ${\theta}^{t}$ = ${\theta^{t}_{1}}$, $\theta^{t}_{2}$,…, ${\theta^{t}_{n}}$,…., ${\theta^{t}_{N}}$}, where ${\theta^{t}_{n}}$, at time t, denotes the parameters of the $n^{th}$ trainable layer out of a total of N.
-3.2 Analysis of Adaptive Learning Rate Schemes for CTTA
-In this subsection, we provide a detailed view and analysis of surgical fine-tuning [21], AutoLR [35], and LAW [33].
-This is essential to understand the design principles and reasons for having an adaptive learning rate scheme for CTTA, which also includes the motivation of our work.
-To preserve the information during the source pre-training phase while fine-tuning on a target task of a different distribution, surgical fine-tuning [21] empirically show that heuristically fine-tuning only the first convolutional block results in a better performance over full fine-tuning.
+CTTA aims to adapt a discriminative source model, parameterized as $\theta^{s}$ and trained on the source dataset $(\mathcal{X}^{\text{source}},\mathcal{Y}^{\text{source}})$, to multiple incoming tasks $\mathcal{T}_{i}=\{x_{k}\}_{k=1}^{K}$, at test-time.
+Each $\mathcal{T}_{i}$ signifies the $i^{th}$ task with $K$ batches, across varying data distributions.
+At time step t=0, the model is initialized to ${\theta^{s}}$.
+It is then gradually adapted to each incoming batch $x_{k}$ of $\mathcal{T}_{i}$ in an online manner, where the model parameters are updated to ${\theta^{t}}$.
+The source dataset is unavailable in this adaptation process due to privacy and storage constraints.
+With the continuum of tasks, at each $t$, the complete model parameters $\theta^{t}$ consist of a deep tensor with $N$ layers, expressed as $\theta^{t}=\{\theta^{t}_{1},\ldots,\theta^{t}_{N}\}$.
+Each layer $\theta^{t}_{n}$ can represented by $\theta^{t}_{n}=\{\theta^{t}_{1,n},\ldots,\theta^{t}_{j,n},\ldots\}$, where $\theta_{j,n}^{t}$ is an arbitrary parameter of layer $\theta_{n}^{t}$.
+3.2 Core Motivations
+We provide a detailed view of surgical fine-tuning [15], AutoLR [26], and LAW [24].
+It is critical to understand the design principles and reasons for having an adaptive LR scheme for CTTA, which motivates this work.
+To preserve the information during the source pre-training phase while fine-tuning on a target task of a different distribution like an image corruption, surgical fine-tuning empirically shows that heuristically fine-tuning only the first convolutional block results in better performance over full fine-tuning.
 Even though it is a desirable outcome for efficient fine-tuning, in this work, we close the gap by automatically selecting the layers that would adapt well to the target task.
-Inspired by [47, 49], AutoLR [35] claims that the weight changes across layers should be in increasing order, where the weight change $\delta w_{n}$ in the $n^{th}$ layer with weights $w_{n}$, at epoch e, is defined as,
-$\delta w_{n}=\|w_{n}^{e}\ -w_{n}^{(e-1)}\|$ (1)
-They hypothesize that weight changes in the initial layers should be small because they extract general features whereas the deeper layers extract task-specific features.
-A constrained adaptive LR optimization scheme is set up to force the weight variations to increase with the layers without studying the importance of each layer during adaptation.
-However, the updated LRs, for each layer, do not deviate too much and remain constant throughout.
-As the first adaptive LR work for CTTA, LAW [33] hypothesizes that identifying and controlling the adaptation rate of the different layers is useful [21].
-To understand the loss landscape and layer importance, the authors compute the gradient from the log-likelihood as,
-$s(\theta_{n}^{t};x_{k})=\nabla_{\theta_{n}^{t}}\log(p_{\theta^{t}}(x_{k}))$ (2)
-where $x_{k}$ and $\theta_{n}^{t}$ refer to the current test batch and parameters of layer n at time t.
-Then, the Fisher Information Matrix (FIM) [36], as a domain-shift indicator, is computed as a first-order approximation of the Hessian matrix.
-As it can be seen, the layer-wise importance (Eq.
-2) completely relies on the pseudo-labels to compute the log-likelihood, which is highly unreliable and uncertain.
+Inspired by [40, 43], AutoLR, proposed for supervised model fine-tuning on a downstream task, claims that the weight changes across layers should be in an increasing order, where the weight change $\delta w_{n}$ in the $n^{th}$ layer with weights $w_{n}$ at training epoch e is $\delta w_{n}=\|w_{n}^{e}\ -w_{n}^{(e-1)}\|$.
+They hypothesize that initial layers, which extract general features, should have minimal $\delta w_{n}$, while deeper layers extract task-specific features.
+An adaptive LR optimization scheme constrains weight variations to increase across layers, without assessing each layer’s importance during adaptation.
+However, the LRs for each layer show minimal deviation and remain consistent throughout the training process.
+As the first adaptive LR work for CTTA, LAW hypothesizes that identifying and controlling the adaptation rate of the different layers is useful [15].
+To understand the loss landscape and layer importance, the authors compute the gradient from the log-likelihood as $s(\theta_{j,n}^{t};x_{k})=\nabla_{\theta_{j,n}^{t}}\log(p(x_{k}))$, where $p(x_{k})$ is the predicted output.
+Then, the FIM [27], as a domain-shift indicator, is computed as a first-order approximation of the Hessian matrix.
+This layer-wise importance scheme completely relies on the pseudo-labels to compute the log-likelihood and is highly unreliable and uncertain.
 Due to the distribution shift, the label quality would begin to decrease for a continuum of tasks.
-Furthermore, the authors propose to capture "domain-level" FIM ($\hat{F_{n}^{t}}$), for each layer during the adaptation, by a direct accumulation of the current layer-wise FIM (${F_{n}^{t}}$) as,
-$\hat{F}_{n}^{t}=\hat{F}_{n}^{t-1}+F_{n}^{t}$ (3)
-Based on Eq.
-3, this means that LAW [33] puts equal emphasis on $\hat{F}_{n}^{t}$ and ${F_{n}^{t}}$, for each layer.
-An issue arises when ${F}_{n}^{t}$ of the current batch is small but gets outweighed by a "relatively" larger $\hat{F}_{n}^{t}$.
+The authors also propose to capture “domain-level" FIM ($\hat{F_{n}^{t}}$), for each layer, by a direct accumulation of the current layer-wise FIM (${F_{n}^{t}}$) as, $\hat{F}_{n}^{t}=\hat{F}_{n}^{t-1}+F_{n}^{t}$.
+This means that LAW puts equal emphasis on $\hat{F}_{n}^{t}$ and ${F_{n}^{t}}$, for each layer.
+An issue arises when ${F}_{n}^{t}$ of the current batch is small but gets outweighed by a “relatively" larger $\hat{F}_{n}^{t}$.
 This raises questions on the amount of adaptation required by the $n^{th}$ layer for batch $x_{k}$ as the FIMs dynamically change for each test batch.
-This leads to an imbalance between the accumulated FIMs $\hat{F}_{n}^{t}$ and the current FIM ${F_{n}^{t}}$.
+This leads to a possible imbalance between the accumulated FIMs $\hat{F}_{n}^{t}$ and the current FIM ${F_{n}^{t}}$.
 Drawing from these primary observations, our method focuses on addressing the following: 1) Based on the prediction uncertainty of the model, we automatically select a subset of layers that would require further adaptation to a task, without relying on the pseudo-labels.
-2) As an indicator of the domain shift, we leverage the idea of parameter sensitivity of these layers only and adaptively alter their learning rates.
-We freeze the layers that are not selected.
+2) As a domain shift indicator, we leverage the idea of parameter sensitivity of these layers only and adaptively alter their learning rates.
+We freeze the parameters of the unselected layers.
 4 Proposed Method
@@
 The parameter sensitivities of these layers, as an indicator of domain shift meters, are computed to update their learning rates (Eq.
-11).
+7).
 Finally, with the optimization objective in Eq.
-16, we update the model with the adjusted learning rates of the parameters.
+11, we update the model with the adjusted learning rates of the parameters.
 In this section, we introduce and formalize our proposed method for CTTA to address the key limitations of the works discussed above. Fig. 1 diagrammatically illustrates our entire approach.
 4.1 Layer Selection by Quantifying Prediction Uncertainty
-CTTA maximizes model performance in the current domain and is vital to assess prediction certainty amidst domain shifts.
+CTTA maximizes model performance in the current domain and so it is vital to assess prediction certainty amidst rapid domain shifts.
 In the literature on continual learning and TTA, several works utilize FIM [13, 29, 24].
-However, the unavailability of true labels, with an unknown distribution, motivates us to capture the prediction uncertainty.
-Prior literature proposed methods for uncertainty from the classifier output or feature space [12, 14, 24, 25, 20].
+However, the unavailability of true labels, with an unknown distribution, pushes us to capture the model prediction uncertainty and adapt selected layers, automatically chosen, to a batch of test data, unlike surgical fine-tuning [15].
 Motivated by [11], which proposed an uncertainty scoring function from the gradient space, we compute the gradients of each layer.
-In essence, we backpropagate the KL divergence (KL) between the softmax probabilities and a uniform distribution u = [$\frac{1}{C}$, $\frac{1}{C}$, …., $\frac{1}{C}$] $\in$ $\mathbb{R}^{C}$, where C is the total number of classes.
+In essence, we backpropagate the KL divergence between the softmax probabilities and a uniform distribution u = [$\frac{1}{C}$, …, $\frac{1}{C}$] $\in$ $\mathbb{R}^{C}$, where $C$ is the total number of classes of the current task.
 This is a measurement of how far the distribution of the model’s prediction is from a uniform distribution.
-Intuitively, due to the distribution shift, the KL divergence should be smaller.
-Let h denote the classifier of the model parameterized by ${\theta^{t}}$.
-We compute the loss, denoting it as $\mathcal{L}(\theta^{t})$ at time stamp t, as follows,
-$\mathcal{L}(\theta^{t})=\text{KL}(\text{softmax}(\hat{h}(x_{k}))||\text{{u}})$ (4)
-where $\hat{h}(x_{k})$ is the classifier’s output logits ($h(x_{k})$) smoothed by a temperature coefficient T i.e.
-$\hat{h}(x_{k})$ = $h(x_{k})$/T.
-Our intuition is that batches within the same task are of the same distribution and so, smoothing the logits with large values of T will lead to them being indistinguishable.
-With model parameters $\theta^{t}$, the gradients (${{\nabla\mathcal{L}({\theta}^{t})}}$ = $\frac{\partial\mathcal{L}(\theta^{t})}{\partial{\theta}^{t}}$) can be simplified to,
-${{\nabla\mathcal{L}({\theta}^{t})}}=\frac{1}{C}\sum_{j=1}^{C}\frac{\partial\mathcal{L_{\text{CE}}}(\hat{h}(x_{k}),j)}{\partial\theta^{t}}$ (5)
-On the whole, the gradients boil down to the averaged cross-entropy loss between the smoothed output logits and a uniform label (Eq.
-5).
-In test-time adaptation, since ground-truth labels are not present, the gradients are more indicative of familiarity of the model with the batch of data.
+Intuitively, the KL divergence will be small with a large distribution shift.
+Let $h(\cdot)$ denote the classifier of the model.
+The loss $\mathcal{L}(\theta^{t})$, to compute the gradients, is computed as follows,
+$\mathcal{L}(\theta^{t})=\text{KL}(\text{softmax}(h(x_{k})/T)||\text{{u}})$ (1)
+where the classifier’s output logits $h(x_{k})$ is smoothed by a temperature $T$.
+Let $\hat{h}(x_{k})$ = $h(x_{k})/T$.
+Our intuition is that since batches within the same task are of the same distribution, smoothing the logits with large values of $T$ will make them indistinguishable.
+With model parameters $\theta^{t}$, the gradients (${{\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta}^{t})}}$ = $\frac{\partial\mathcal{L}(\theta^{t})}{\partial\theta_{j,n}^{t}}$) can be simplified to,
+${{\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta}^{t})}}=\frac{1}{C}\sum_{c=1}^{C}\frac{\partial\mathcal{L_{\text{CE}}}(\hat{h}(x_{k}),c)}{\partial\theta_{j,n}^{t}}$ (2)
+Overall, the gradients boil down to the averaged cross-entropy loss between the smoothed output logits and a uniform label (Eq.
+2).
+In CTTA, since ground-truth labels are absent, the gradients are more indicative of familiarity of the model with the batch of data.
 For batches within the same task, that share the same distribution, the vector norm of Eq.
@@
 However, if the norm is smaller, this would mean that the softmax outputs are closer to u.
-Our key idea is to select all the parameters of layers that have the norm of gradients to be less than or equal to a threshold $\eta$.
+Our key idea is to select all the parameters of layers with the gradient norms to be less than or equal to a threshold $\eta$.
 Layers with small gradient norms would require more adaptation for the model to make a confident prediction.
-Since we freeze the remaining layers, this also helps to alleviate catastrophic forgetting to a large extent by preserving the prior knowledge.
-And so the scoring function, for each layer, is defined as,
-$\mathcal{Z}_{\theta_{n}}=\|{{\nabla\mathcal{L}({\theta}_{n}^{t})}}\|_{1}$ (6)
-As mentioned earlier, $\mathcal{Z}_{\theta_{n}}$leq $\eta$ allows the model to adaptively determine which specific layers need further adaptation.
-We freeze the layers with $\mathcal{Z}_{\theta_{n}}$ $>$ $\eta$ i.e.
-set their learning rate to 0.
-This way, in contrast to surgical fine-tuning [21], we automatically select a subset of layers for adaptation to a batch of data.
-4.2 Parameter Sensitivity as an Indicator of Domain Shift
+We freeze the unselected layers which also helps to alleviate catastrophic forgetting to a large extent by preserving the prior source knowledge.
+So, the scoring function, for each layer’s parameters, is defined as $\mathcal{Z}_{\theta_{j,n}^{t}}=\|{\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta}^{t})}\|_{1}$.
+As mentioned earlier, $\mathcal{Z}_{\theta_{j,n}^{t}}$leq $\eta$ allows the model to adaptively determine which specific layers need further adaptation.
+We freeze the layers with $\mathcal{Z}_{\theta_{j,n}^{t}}$ $>$ $\eta$ i.e., set their LR to 0.
+4.2 Parameter Sensitivity as a Domain Shift Indicator
 New studies have found that model parameters demonstrate levels of criticality [2, 44] across different positions.
 Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors [22, 25].
-Inspired by this observation, we claim that parameters have different contributions during test-time adaptation.
-From Sec.
-4.1, we adaptively select layers that are uncertain in their predictions.
-For parameters of these layers only, we compute their sensitivity i.e.
-an approximation of an error when it is removed.
-In our work, motivated from [19, 30, 29], we compute the parameter sensitivity of a parameter $\theta_{n}^{t}$ based on the loss (Eq. 4) as follows,
-$S_{n}^{t}=\mathcal{L}({\theta}^{t})-\mathcal{L}({\theta}^{t}|\theta_{n}^{t}=0)$ (7)
-Precisely, $S_{n}^{t}$ approximates the change in the loss landscape at $\theta^{t}$ when $\theta_{n}^{t}$ is removed.
-Sensitivity, computed as a "weighting factor" of model weights [30, 29, 23], determines the level of adaptation needed for model parameters.
-We utilize the gradients of parameters from the selected layers, computed by Eq. 5 to compactly derive Eq. 7 from the first-order Taylor series expansion of $\mathcal{L}(\theta^{t})$ as,
-$S_{n}^{t}=|\theta_{n}^{t}\nabla\mathcal{L}({\theta_{n}}^{t})|$ (8)
-Our rationale is as follows: parameters with low sensitivity are pushed to have higher learning rates owing to more adaptation.
-However, $S_{n}^{t}$ of $\theta_{n}^{t}$ can be highly uncertain and would have a large variation.
-We introduce a method to capture this "domain-level" uncertainty by computing a weighted moving average of the sensitivity $S_{n}^{t}$.
+Inspired by this observation, we assert that parameters contribute differently during CTTA.
+From the previous stage, we automatically select the layers based on the prediction uncertainty.
+For parameters of these selected layers only, we compute their sensitivity i.e., an approximation of an error when it is removed.
+Motivated by [14, 21, 20], we compute the sensitivity of a parameter $\theta_{j,n}^{t}$ based on the loss (Eq.
+1) as follows,
+$S_{j,n}^{t}=\mathcal{L}({\theta}^{t})-\mathcal{L}({\theta}^{t}|\theta_{j,n}^{t}=0)$ (3)
+Precisely, $S_{j,n}^{t}$ approximates the change in the loss landscape at $\theta^{t}$ when $\theta_{j,n}^{t}$ is removed.
+Sensitivity, computed as a “weighting factor" of model weights [21, 20, 16], determines the level of adaptation needed for model parameters.
+We utilize the gradients of parameters from the selected layers, computed by Eq.
+2, to compactly derive Eq.
+3 from the first-order Taylor series expansion of $\mathcal{L}(\theta^{t})$ [lee2018snip] to,
+$S_{j,n}^{t}=|\theta_{j,n}^{t}\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta^{t}})|$ (4)
+Our rationale is as follows: parameters with low sensitivity are pushed to have higher LRs owing to more adaptation.
+However, $S_{j,n}^{t}$ of $\theta_{j,n}^{t}$ can be highly uncertain and would have a large variation.
+We introduce a method to capture this “domain-level" uncertainty by computing a weighted moving average of the sensitivity $S_{j,n}^{t}$.
 This is formulated as,
-$\hat{S}_{n}^{t}=\alpha S_{n}^{t}+(1-\alpha)\hat{S}_{n}^{(t-1)}$ (9)
-where $\hat{S}_{n}^{t}$ denotes the "domain-level" sensitivity and $\alpha$in [0,1] is a smoothing factor.
-Specifically, $\alpha$ decides the contribution of the current sensitivity $S_{n}^{t}$ to $\hat{S}_{n}^{t}$.
-Since CTTA happens at batch level, it is critical to achieve a good balance between $S_{n}^{t}$ and $\hat{S}_{n}^{t}$, with a slightly larger focus on $S_{n}^{t}$.
-As it can be seen, we do not directly accumulate the "domain-level" information, unlike LAW [33].
-In addition, due to large distribution shifts happening because of the continuum of tasks, larger deviations of $S_{n}^{t}$ from $\hat{S}_{n}^{t}$ are bound to happen.
-To quantify these statistics, an absolute difference between $\hat{S}_{n}^{t}$ and $S_{n}^{t}$ can be computed as,
-$\hat{D}_{n}^{t}=|S_{n}^{t}-\hat{S}_{n}^{t}|$ (10)
+$\hat{S}_{j,n}^{t}=\alpha S_{j,n}^{t}+(1-\alpha)\hat{S}_{j,n}^{(t-1)}$ (5)
+where $\hat{S}_{j,n}^{t}$ denotes the “domain-level" sensitivity, computed on-the-fly, and $\alpha$in [0,1] is a smoothing factor.
+Specifically, $\alpha$ decides the contribution of the current sensitivity $S_{j,n}^{t}$ to $\hat{S}_{j,n}^{t}$.
+Since CTTA happens at the batch level, it is critical to achieve a good balance between $S_{j,n}^{t}$ and $\hat{S}_{j,n}^{t}$, with a slightly larger focus on $S_{j,n}^{t}$.
+As can be seen, we do not directly accumulate the “domain-level" information, unlike LAW.
+Due to large distribution shifts happening because of the continuum of tasks, larger deviations of $S_{n}^{t}$ from $\hat{S}_{n}^{t}$ are bound to happen.
+We quantify this statistic as an absolute difference,
+$\hat{D}_{j,n}^{t}=|S_{j,n}^{t}-\hat{S}_{j,n}^{t}|$ (6)
 We emphasize that since Eq.
-9 leverages sensitivity from previously computed batches and domains, $\hat{D}_{n}^{t}$ would act as a good indicator of how uncertain $S_{n}^{t}$ is to the previous batches and domains, for $\theta_{n}$.
+5 leverages sensitivity from previously computed batches and domains, $\hat{D}_{j,n}^{t}$ would act as a good indicator of how uncertain $S_{j,n}^{t}$ is to the previous batches and domains i.e.
+$\hat{S}_{j,n}^{t}$, for $\theta_{j,n}^{t}$.
 This would take care of error accumulation too.
-To automatically alter the learning rates of these parameters, the learning rate "importance" ($i_{n}^{t}$) is calculated as follows,
-$i_{n}^{t}=(\hat{D}_{n}^{t})/(\hat{S}_{n}^{t})$ (11)
-Then, the final learning rate to update $\theta_{n}^{t}$ would be
-$\hat{\kappa}_{n}^{t}=\kappa*i_{n}^{t}$ (12)
-where $\kappa$ is the base fixed learning rate.
+To automatically alter the LRs of these parameters, the “importance" ($i_{j,n}^{t}$) is calculated as,
+$i_{j,n}^{t}=\hat{D}_{j,n}^{t}/\hat{S}_{j,n}^{t}$ (7)
+Eq. 6 enforces $\hat{D}_{j,n}^{t}$$\geq$0 which, in turn, makes $i_{j,n}^{t}$$\geq$0. And so, the final learning rate to update $\theta_{j,n}^{t}$ would be,
+$\hat{\kappa}_{j,n}^{t}=\kappa*i_{j,n}^{t}$ (8)
+where $\kappa$ is the base fixed LR.
 The intuition behind the formulation of Eq.
-11 is as follows - smaller values of $\hat{S}_{n}^{t}$ indicate lower sensitivity and hence larger parameter updates are required via promoting the learning rate.
-At the same time, large uncertainties by $\hat{D}_{n}^{t}$ would also require larger updates.
+7 is - smaller values of $\hat{S}_{j,n}^{t}$ indicate lower sensitivity, and hence larger parameter updates are required via promoting the LR.
+At the same time, large uncertainties by $\hat{D}_{j,n}^{t}$ would also require larger updates.
 To avoid any overflow/underflow, we add $\epsilon$ (0$<$$\epsilon$$\ll$1) to the numerator and the denominator in Eq.
-11.
+7.
 4.3 Optimization Objective
-Though our method is robust, it is necessary to avoid miscalibrated labels. TENT [44] minimized the Shannon entropy on the output logits, defined as,
-$\mathcal{H}(\hat{y})=-\sum_{c}p(y_{c})log(p(y_{c}))$ (13)
-where $p(y_{c})$ refers to the softmax probability of class c i.e.
-$p(y_{c})$ = softmax($h(x_{k})$).
-Following previous CTTA work [41], we use entropy minimization as the main loss objective for adaptation where Eq.
-13 is minimized for samples below a certain threshold.
-$\mathcal{L}^{entropy}=\mathds{1}_{\mathcal{H}(\hat{y})\leq\mathcal{H}_{0}}\mathcal{H}(\hat{y})$ (14)
-where $\mathds{1}$ is the indicator function and $\mathcal{H}_{0}$ is a fixed threshold set to $0.4*lnC$, consistently following [41].
-To be fair with prior works [45, 33], we use the consistency loss between the model predictions on $x_{k}$ and its augmented version $\hat{x}_{k}$ to regularize the model as follows,
-$\mathcal{L}^{const}=-\sum_{c}p(y_{c})log(p(\hat{y}_{c}))$ (15)
-where, $\hat{y}_{c}$ = $h(\hat{x}_{k})$. Our overall optimization objective is as follows,
+Though our method is robust, it is necessary to avoid miscalibrated labels.
+TENT [36] minimized the entropy on the output logits, l, following $\mathcal{H}(l)=-\sum_{c}p(l_{c})\log(p(l_{c}))$, where $p(l_{c})$ is the probability for class c.
+Following EcoTTA [33], we use entropy minimization as the objective for adaptation where $\mathcal{H}(l)$ is minimized for samples below a certain fixed threshold $\mathcal{H}_{0}$ as,
+$\mathcal{L}^{entropy}=\mathbf{1}_{\mathcal{H}(l)\leq\mathcal{H}_{0}}\mathcal{H}(l)$ (9)
+$\mathcal{H}_{0}$ is set to $0.4\times\log C$ following EcoTTA.
+For a fair comparison with prior works [37, 24], we apply a consistency loss between l and $\hat{l}$, computed as the output logits from the augmented version of $x_{k}$, to regularize the model as follows,
+$\mathcal{L}^{const}=-\sum_{c}p(l_{c})\log(p(\hat{l}_{c}))$ (10)
+where, $p(\hat{l}_{c})$ is the probability for class c. Our overall optimization objective is as follows,
 $\mathcal{L}^{total}=\mathcal{L}^{entropy}+\lambda\mathcal{L}^{const}$ (11)
-where $\lambda$ indicates a regularisation coefficient to balance the consistency loss.
-In summary, upon calculating the parameter-specific learning rates (Eq.
-12), we optimize the continually fine-tuned model to minimize Eq.
-16.
+where $\lambda$ is a regularisation coefficient. In summary, upon calculating the parameter-specific LRs (Eq. 8), we optimize the continually fine-tuned model to minimize Eq. 11.
 5 Experiments
@@
 In the following subsections, we describe the implementation details of both the challenging adaptation settings.
-For additional results and analysis, please refer to the Supplementary Material.
 5.1 Experimental Settings and Baselines
 Datasets and tasks.
-Following the standard benchmarks in CTTA, set by CoTTA [45], we evaluate our proposed method on datasets - CIFAR-10C, CIFAR-100C, and ImageNet-C, based on image corruption schemes as set in [11].
-Each dataset contains a set of 15 corruption styles (e.g.
-gaussian noise, shot noise, impulse noise, defocus blur) with 5 severity levels.
-In line with established practices outlined in [45, 6], we report the mean classification error as the evaluation metric.
+Following the standard benchmarks set by CoTTA [37], we evaluate our proposed method on CIFAR-10C, CIFAR-100C, and ImageNet-C, based on image corruption schemes as set in [9].
+Each dataset contains a set of 15 corruption styles as tasks (e.g.
+gaussian noise, shot noise, impulse noise, $\ldots$) with 5 severity levels.
+We follow CoTTA and apply the same test-time augmentations, like color jitter, gaussian blur, gaussian noise, random flips, and random affine, to the input $x_{k}$.
+As the evaluation metric, we report the mean classification error.
 CTTA implementation details.
-We evaluate our approach using frameworks from prior state-of-the-art methods [44, 45, 3, 41, 37, 33].
-To maintain fairness regarding the usage of pre-trained vision models, we employ WideResNet-28 [48] for CIFAR-10C, ResNeXt-29 [46] for CIFAR-100C, and ResNet-50 [10] for ImageNet-C, all available on RobustBench [5].
-We optimize using an Adam optimizer [17], setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C.
+We evaluate our approach following prior state-of-the-art methods [36, 37, 3, 33, 28, 24].
+To maintain fairness regarding the usage of pre-trained source models, we employ WideResNet-28 [42] for CIFAR-10C, ResNeXt-29 [39] for CIFAR-100C, and ResNet-50 [8] for ImageNet-C, all available on RobustBench [4].
+We optimize using an Adam optimizer, setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C.
 For balancing parameter sensitivities, we set $\alpha$ to 0.5, 0.9, and 0.5 respectively, with temperature coefficients T set to 50, 100, and 1000 respectively.
 We set $\eta$ to 1, 0.5, and 0.3, and $\lambda$ to 0.01 throughout.
-Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA [45], and results are reported for the highest severity level of 5 for each task.
+Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA, and results are only reported for the highest severity level of 5 for each task.
 GTTA implementation details.
-In the standard CTTA setup, each task receives images with the highest severity.
-However, in GTTA, following [45, 41, 33], for each task the severity levels are gradually changed as …3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5$\rightarrow$4$\rightarrow$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3… We maintain consistency with the experimental details outlined in the CTTA setting.
+In the standard CTTA setup, each task only receives images with the highest severity.
+However, in GTTA, following CoTTA, for each task, the severity levels are gradually changed as $\ldots$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5$\rightarrow$4$\rightarrow$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\ldots$.
+We adhere to the details specified in the CTTA setting.
 Baselines.
-We compare our work against several CTTA baselines, including Source, BN Stats Adapt [37], TENT Continual [44], CoTTA [45], EcoTTA [41], Adacontrast [3], and LAW [33].
-In addition, we also contrast our work against surgical fine-tuning [21].
+We compare against several CTTA baselines, including BN Stats Adapt (BN-1) [28], TENT Continual [36], CoTTA [37], EcoTTA [33], AdaContrast [3], DSS [38], and LAW [24].
+We also contrast our work against surgical fine-tuning [15].
 For EcoTTA, we report the results on CIFAR-10C and ImageNet-C only since they do not use ResNeXt-29 [39] for CIFAR-100C.
-Following the findings of surgical fine-tuning [21], we continually fine-tune only the first convolutional block of the respective models, denoted as Surgical fine-tuning (continual).
-We have mentioned the training details in the Supplementary.
-For the reported results of LAW [33], we reproduce their results strictly following their official implementation  [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
-Table 1: Mean errors (in %) on CIFAR-10C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5.
-We report the error of a single update (online) only using WideResNet-28 [48] for all the reported baselines.
+Following the findings of surgical fine-tuning [15], we continually fine-tune only the first convolutional block of the respective models, denoted as “Surgical fine-tuning (continual)", using an SGD optimizer with fixed learning rates of 1e-3 for the CIFAR datasets and 5e-4 for ImageNet-C.
+For the reported results of LAW, we reproduce their results strictly following their official implementation [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
 Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
@@
 Source 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.5 30.3 43.5
-BN Stats Adapt [37] 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
+BN-1 [28] 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
 TENT Continual [36] 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
@@
 fine-tuning [15]
-27.71 24.14 33.79 12.18 33.07 13.09 10.97 15.94 15.61 13.8 7.84 10.77 21.17 16.62 22.88 18.64
+27.7 24.1 33.8 12.2 33.1 13.1 10.9 15.9 15.6 13.8 7.8 10.8 21.2 16.6 22.9 18.6
+DSS [38]
+24.1 21.3 25.4 11.7 26.9 12.2 10.5 14.5 14.1 12.5 7.8 10.8 18.0 13.1 17.3 16.0
 LAW [24]
-24.71 18.91 25.56 12.96 26.78 15.09 11.8 15.17 14.78 15.89 10.13 13.89 19.41 14.74 18.35 17.21
+24.7 18.9 25.5 12.9 26.7 15.0 11.8 15.1 14.7 15.9 10.1 13.8 19.4 14.7 18.3 17.2
 Ours
-25.89 18.1 22.71 12.38 25.3 13.18 10.79 13.5 13.15 12.21 8.53 11.85 17.98 12.02 15.47 15.54
-Table 2: Mean errors (in %) on CIFAR-100C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5.
-We report the error of a single update (online) only using ResNeXt-29 [46] for all the reported baselines.
+25.8 18.1 22.7 12.3 25.3 13.1 10.7 13.5 13.1 12.2 8.5 11.8 17.9 12.0 15.4 15.5
+Table 1: Mean errors (%) on CIFAR-10C - CTTA mean errors of the 15 corruptions (tasks) at a severity level of 5.
 Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
@@
 73.0 68 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
-BN Stats Adapt [37]
+BN-1 [28]
 42.1 40.7 42.7 27.6 41.9 29.7 27.9 34.9 35.0 41.5 26.5 30.3 35.7 32.9 41.2 35.4
@@
 fine-tuning [15]
-40.75 36.36 37.89 25.53 39.02 27.65 25.27 32.26 30.45 37.35 23.78 26.82 32.37 27.64 37.20 32.02
+40.7 36.3 37.8 25.5 39.0 27.6 25.2 32.2 30.4 37.3 23.7 26.8 32.3 27.6 37.2 32.0
+DSS[38]
+39.7 36.0 37.2 26.3 35.6 27.5 25.1 31.4 30.0 37.8 24.2 26.0 30.0 26.3 31.1 30.9
 LAW [24]
-41.06 36.73 38.38 25.67 37.02 27.82 25.22 30.79 30.04 37.35 24.41 27.69 31.25 27.84 34.92 31.75
+41.0 36.7 38.3 25.6 37.0 27.8 25.2 30.7 30.0 37.3 24.4 27.6 31.2 27.8 34.9 31.7
 Ours
-37.33 32.5 34.94 26.20 35.28 27.54 24.64 28.85 29.15 34.14 23.52 26.99 31.15 26.56 34.1 30.19
+37.3 32.5 34.9 26.2 35.3 27.5 24.6 28.8 29.1 34.1 23.5 26.9 31.2 26.6 34.1 30.1
+Table 2: Mean errors (%) on CIFAR-100C - CTTA mean errors of the 15 corruptions (tasks) at a severity level of 5.
 5.2 CTTA Classification Results
 Results on CIFAR-10C and CIFAR-100C.
-We contrast our proposed method against several baselines including surgical fine-tuning [21].
-Tables 1 and 2 report the CTTA mean classification errors across all the domains, for each of the methods.
+We contrast our method against all the baselines as in Tables 1 and 2.
 The results reveal that our method performs the best across all training setups and baselines for both datasets.
-In particular, we achieve large improvements of 4.86% and 5.16% for CIFAR-10C and 5.21% and 30.71% for CIFAR-100C, over BN Stats Adapt[3] and TENT Continual[44] respectively.
-Concerning recent work LAW [33] that uses an adaptive LR scheme, our method achieves an improvement of 1.67% and 1.56% on CIFAR-10C and CIFAR-100C, respectively.
-It’s noteworthy that TENT Continual initially excels in the early tasks but experiences error accumulation over time, leading to an increase in mean error.
-In the study by CoTTA [45], despite achieving good performance, the approach’s reliance on full model updates renders it computationally expensive and unsuitable for deployment.
-LAW [33] uses the uncertain pseudo labels, of each batch, to calculate the domain shift approximated via the FIM computation.
+We achieve large improvements of 4.9% and 5.2% for CIFAR-10C and 5.3% and 30.8% for CIFAR-100C, over BN-1 and TENT Continual respectively.
+Concerning recent works of LAW and DSS, our method improves 1.7% and 0.5% on CIFAR-10C and 1.6% and 0.8% on CIFAR-100C, respectively.
+It is noteworthy that TENT Continual initially excels in the early tasks but experiences error accumulation with time, leading to an increase in mean error.
+In the study by CoTTA, the method’s reliance on full model updates renders it computationally expensive and unsuitable for deployment.
+An interesting observation is that with a drastic reduction in the number of trainable parameters, as in “Surgical fine-tuning (continual)", the mean errors are better than most CTTA benchmarks, if not the best.
+LAW uses the uncertain pseudo labels to calculate the domain shift approximated via the FIM.
 The minor error that is introduced for each batch begins to propagate across all the domains.
-In addition, the imbalance caused by the "domain-level" FIM and the current batch’s FIM, results in a higher error rate for the initial tasks.
-In comparison, the computation of the model prediction uncertainty to select layers is useful in our work (Eq.
-4).
-Another interesting observation is that with a drastic reduction in the number of trainable parameters, as in Surgical fine-tuning (continual), the mean errors in Table 2 is better than most CTTA benchmarks, if not the best.
-This affirms the findings in [21] that concentrated adaptation, for image corruptions, in the initial layers is sufficient.
-Results on Imagenet-C In Table 3, we report the CTTA results on the ImageNet-C dataset, with ResNet-50 pre-trained on ImageNet.
-As shown, our method outperforms all the previous CTTA baselines and Surgical fine-tuning (continual).
-We observe an improvement of 1% over LAW [33].
-In addition, we achieve 2.31%, 5.38%, and 30.71% improvements over CoTTA [45], Adacontrast [3], and TENT [44], respectively.
-Table 3: Mean errors (in %) on ImageNet-C - We report the CTTA mean errors of the 15 corruptions at a severity level of 5.
-We report the error of a single update (online) only using ResNet-50 [10] for all the reported baselines.
+In addition, the imbalance caused by the “domain-level" FIM and the current batch’s FIM, results in a higher error rate for the initial tasks.
+In our method, the computation of the model prediction uncertainty to automatically select layers is useful (Eq.
+1).
+This affirms the findings in [15] that concentrated adaptation, for image corruptions, in the initial layers suffices.
+Results on Imagenet-C In Table 3, we report the CTTA results on the ImageNet-C dataset.
+As shown, our method outperforms all the previous CTTA baselines and “Surgical fine-tuning (continual)".
+We observe an improvement of 0.9% and 4.5% over LAW and DSS respectively.
+In addition, we achieve 2.6%, 5.4%, and 2.5% improvements over CoTTA, AdaContrast, and TENT Continual, respectively.
 Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
@@
 97.8 97.1 98.2 81.7 89.8 85.2 78.0 83.5 77.1 75.9 41.3 94.5 82.5 79.3 68.6 82.0
-BN Stats Adapt [37]
+BN-1 [28]
 85.0 83.7 85.0 84.7 84.3 73.7 61.2 66.0 68.2 52.1 34.9 82.7 55.9 51.3 59.8 68.6
@@
 fine-tuning [15]
-83.42 79.5 77.28 86.1 82.4 72.28 61.18 63.1 64.76 51.26 35.92 72.9 56.24 46.72 50.2 65.55
+83.4 79.5 77.2 86.1 82.4 72.2 61.1 63.1 64.7 51.2 35.9 72.9 56.2 46.7 50.2 65.5
+DSS [38]
+84.6 80.4 78.7 83.9 79.8 74.9 62.9 62.8 62.9 49.7 37.4 71.0 49.5 42.9 48.2 64.6
 LAW [24]
-81.86 75.04 72.12 77.42 73.22 63.96 54.6 57.92 61.06 50.1 36.36 68.66 49.08 46.22 49.12 61.12
+81.8 75.0 72.1 77.4 73.2 63.9 54.6 57.9 61.0 50.1 36.3 68.6 49.0 46.2 49.1 61.0
 Ours
-81.06 73.32 70.88 77.00 71.88 62.34 53.94 56.7 60.76 50.36 36.26 65.96 48.08 45.28 48.0 60.12
-Table 4: Mean errors (in %) on CIFAR-10C, CIFAR-100C, and ImageNet-C - We report the gradual TTA (GTTA) mean errors averaged across all the severity levels (1normal-$\rightarrow$2normal-$\rightarrow$3normal-$\rightarrow$4normal-$\rightarrow$5).
+81.1 73.3 70.8 77.0 71.8 62.3 53.9 56.7 60.8 50.4 36.2 65.9 48.0 45.2 48.0 60.1
+Table 3: Mean errors (%) on ImageNet-C - CTTA mean errors of the 15 corruptions (tasks) at a severity level of 5.
 Dataset Source
-BN Stats
+BN-1
 Adapt
@@
 LAW Ours
-CIFAR-10C 24.7 13.7 20.4 10.9 12.1 11.33 10.87 9.56
-CIFAR-100C 33.6 29.9 74.8 26.3 33.0 25.31 27.19 26.11
-ImageNet-C 58.4 48.3 46.4 38.8 66.3 41.36 40.52 38.49
+CIFAR-10C 24.7 13.7 20.4 10.9 12.1 11.3 10.9 9.5
+CIFAR-100C 33.6 29.9 74.8 26.3 33.0 25.3 27.2 26.1
+ImageNet-C 58.4 48.3 46.4 38.8 66.3 41.4 40.5 38.5
+Table 4: Mean errors (in %) on CIFAR-10C, CIFAR-100C, and ImageNet-C - We report the GTTA mean errors averaged across all the severity levels (1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5).
 5.3 GTTA Classification Results
 Results on CIFAR-10C, CIFAR-100C, and Imagenet-C.
-This subsection presents the gradual test-time adaptation (GTTA) results on CIFAR-10C, CIFAR-100C, and Imagenet-C.
-For each task, the corruption severity of the inputs is gradually increased from levels 1 to 5, as discussed in the implementation details in Sec.
-5.1.
+Here, we present the gradual test-time adaptation (GTTA) results on CIFAR-10C, CIFAR-100C, and Imagenet-C.
+For each task, the corruption severity of the inputs is gradually changed as discussed earlier.
 Table 4 reports the mean classification error for GTTA across all severities.
-We observe that, for all the datasets, our approach outperforms or performs comparably well across all the baselines [37, 44, 45, 3, 33].
-With most parameters frozen during the adaptation on each batch, our method is more robust and inexpensive than CoTTA [45], which also demonstrates competitive improvements in the mean classification error.
+We observe that, for all the datasets, our approach outperforms or performs comparably well across all the reported baselines.
+With most parameters frozen during the adaptation on each batch, our method is more robust and inexpensive than CoTTA, which also demonstrates competitive improvements in the mean classification error.
 5.4 Ablation Studies
@@
 Refer to caption Refer to caption Refer to caption
-Figure 2: Illustrations of the learning rate importances of WideResNet-28 [48], ResNeXt-29 [46], and ResNet-50 [10] during CTTA, for different convolutional blocks/stages/layers.
-[Top row] - Variations across respective datasets for domain glass_blur.
-[Bottom row] - Variations across respective datasets for domain snow.
-Analysis of the learning rate importance.
-The prime objective of our work is to adaptively alter the learning rates of parameters of selected layers based on their sensitivity.
-In Fig.
-2, we illustrate how the learning rate (LR) importance (Eq.
-11) of WideResNet-28 [48], ResNeXt-29 [46], and ResNet-50 [10], averaged for a task, change during adaptation on their respective datasets.
+Figure 2: Illustrations of LR importance $i_{j,n}^{t}$ across different convolutional blocks/stages/layers during CTTA.
+[Top row] - Variations for the domain “glass_blur" across datasets.
+[Bottom row] - Variations for the domain “snow".
+Dataset Method
+The average % of
+adapted parameters/task
+(source parameters)
+CIFAR-10C Ours 4.34 (of 36.4M)
+CoTTA 100 (of 36.4M)
+CIFAR-100C Ours 6.65 (of 6.8M)
+CoTTA 100 (of 6.8M)
+ImageNet-C Ours 17.21 (of 23.5M)
+CoTTA 100 (of 23.5M)
+Table 5: Average % of adapted parameters/task.
+Computational efficiency and Catastrophic forgetting: In our proposed method, we first select a subset of layers from the gradient space (Eq.
+2) and adapt them while strictly freezing the unselected layers.
+In Table 5, we report the average % of adapted model parameters per task with respect to the source model, for our method and CoTTA, across the benchmark datasets.
+Our method is computationally inexpensive and on an average updates 4.34%, 6.65%, and 17.21% of their respective source model parameters.
+This showcases significant computational advantages over CoTTA, which updates 100% of the source parameters.
+Additionally, in LAW, the LRs of all the parameters are adjusted.
+It is worth mentioning that our method alleviates catastrophic forgetting (CF) by only adapting a small subset of model parameters ensuring that much of the prior source knowledge is preserved and CF is largely minimized.
+Dataset Ours LARS Prodigy Adafactor
+CIFAR-10C 15.5 18.9 20.4 73.8
+CIFAR-100C 30.1 34.1 35.5 97.7
+ImageNet-C 60.1 62.0 68.6 99.2
+Table 6: CTTA mean errors (%) using adaptive LR optimizers vs. ours.
+Adaptive LR optimizers vs Ours: Here, we study the effect of directly using adaptive LR optimizers like LARS [41], Prodigy [19], and Adafactor [32], for CTTA.
+We report the results in Table 6.
+The mentioned optimizers do not adjust layer-wise LRs based on parameter sensitivity, failing to capture rapid domain shifts, resulting in poorer performance compared to our method.
+Analysis of LR importance $i_{j,n}^{t}$: The prime objective of our work is to adaptively alter the LRs of parameters of selected layers based on their sensitivity.
+In Figure 2, we illustrate how the LR importance $i_{j,n}^{t}$ (Eq.
+7) of WideResNet-28, ResNeXt-29, and ResNet-50, averaged for a task, change during adaptation on their respective datasets.
 For brevity, we show two domains - glass_blur and snow.
-As we can see, the importance varies across domains and models.
-Across all the models, we notice that generally, the deeper layers have smaller associated LR importances, which means that the initial layers are adapted more to the target domain.
+We observe that the importance varies across domains and models.
+Across all the models, we notice that generally, the deeper layers have smaller associated $i_{j,n}^{t}$, which means that the initial layers are adapted more to the target domain.
 This result completely aligns with the findings of surgical fine-tuning [15].
@@
 This further confirms the necessity of freezing parameters automatically selected based on the target domain for adaptation.
-Nullifying prediction uncertainty.
-In LAW [33], the authors make use of the pseudo-labels to automatically decide on the importance of layers for adaptation.
-In this, we study the effectiveness of such a design for our method i.e.
-we investigate if layer importance computed using pseudo-labels is helpful.
-We compute the gradients from the log-likelihood (Eq.
-2) and estimate the sensitivity of the parameters of the chosen layers.
-Table 5.4 reports the mean classification errors for all the datasets.
+Layer importance via pseudo-labels: In LAW, the authors use the pseudo-labels to decide on the importance of layers for adaptation.
+In this, we study the effectiveness of such a design for our method i.e., we investigate if layer importance computed using pseudo-labels is helpful.
+We compute the gradients from the log-likelihood as $\nabla_{\theta_{j,n}^{t}}\log(p(x_{k}))$ and estimate the sensitivity of the parameters of the chosen layers (Eq.
+4-Eq.
+6), and alter the learning rates (Eq.
+7).
+Table 7 reports the mean classification errors for all the datasets.
 We observe that the gradients computed using the pseudo-labels are an unreliable indicator of prediction uncertainty.
-The scoring function (Eq.
-6) computed using these gradients carries insufficient information.
-We achieve improvements of 2.44%, 3.69%, and 8.5% over CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
+The scoring function $\mathcal{Z}_{\theta_{j,n}^{t}}=\|{{\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta}^{t})}}\|_{1}$ computed using these gradients carries insufficient information.
+We achieve improvements of 2.4%, 3.7%, and 8.5% over CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
 Refer to caption Refer to caption
@@
 Figure 3: Ablation results on smoothing factor $\alpha$ and temperature T for CIFAR-10C, CIFAR-100C, and ImageNet-C.
-Table 5: Prediction uncertainty using pseudo-labels vs our method. Dataset Pseudo-labels Ours CIFAR-10C 17.98 15.54 CIFAR-100C 33.88 30.19 ImageNet-C 68.62 60.12
-Table 6: Ablation results on benchmark datasets with varying $\lambda$.
-$\lambda$ 0 0.01 0.05 0.1 0.5 1.0 CIFAR-10C 16.84 15.54 16.45 17.03 17.65 17.93 CIFAR-100C 30.44 30.19 30.51 30.57 30.63 30.7 ImageNet-C 63.98 60.12 60.38 60.4 60.4 60.42
-Contribution of smoothing factor $\alpha$.
-An important component of our method is the computation of the moving average of sensitivities across data batches, smoothed by a parameter $\alpha$ (Eq.
-9).
+Dataset Pseudo-labels Ours
+CIFAR-10C 17.9 15.5
+CIFAR-100C 33.8 30.1
+ImageNet-C 68.6 60.1
+Table 7: Layer importance using pseudo-labels vs ours.
+$\lambda$ 0 0.01 0.05 0.1 0.5 1.0
+CIFAR-10C 16.8 15.5 16.4 17.0 17.6 17.9
+CIFAR-100C 30.4 30.1 30.5 30.5 30.6 30.7
+ImageNet-C 63.9 60.1 60.3 60.4 60.4 60.4
+Table 8: Ablation results with varying $\lambda$.
+Contribution of smoothing factor $\alpha$: A key component of our method is computing the moving average of sensitivities across data batches, smoothed by parameter $\alpha$ (Eq.
+5).
 Fig.
 3(a) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$.
-We observe that the "domain-level" sensitivity across past batches of data is indeed useful.
-For $\alpha$ = 0 i.e.
-when the current sensitivity of parameters of selected layers is not considered, the mean classification error for the respective datasets drastically spikes up.
+We observe that the “domain-level" sensitivity across past batches of data is indeed useful.
+For $\alpha$ = 0 i.e., when $S_{j,n}^{t}$ of a selected parameter at timestep t is not considered, the respective datasets’ error rate drastically spikes.
 This suggests that the effect of the current sensitivity of selected layers is indeed required.
-For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.67% over LAW [33].
-Similarly, for CIFAR-100C and ImageNet-C, we see improvements of 1.56% and 1% over LAW [33], at $\alpha$ = 0.9 and 0.5 respectively.
-With higher values of $\alpha$, the results do not vary too much and are stable.
-Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{n}^{t}$) and current sensitivity ($S_{n}^{t}$) is critical, as seen in the figure.
+For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.7% over LAW.
+For CIFAR-100C and ImageNet-C, we see improvements of 1.6% and 0.9% over LAW, at $\alpha$ = 0.9 and 0.5 respectively.
+With higher values of $\alpha$, the results do not vary too much.
+Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{j,n}^{t}$) and current sensitivity ($S_{j,n}^{t}$) is needed.
 […]
-Robustness to the regularisation coefficient $\lambda$.
-As an important ablation step, we understand the effect of the parameter $\lambda$ on the final optimization objective, as in Eq.
-16.
+Robustness to regularization $\lambda$: Here, we understand the effect of the parameter $\lambda$ in Eq.
+11.
 We set $\lambda$in {0, 0.01, 0.05, 0.1, 0.5, 1.0}.
-We study the contributions of the consistency loss to entropy minimization.
-Table 5.4 reports the average classification errors on CIFAR-10C, CIFAR-100C, and ImageNet-C, with varying $\lambda$.
-For $\lambda$ = 0.01, we achieve the best results across all the datasets.
-In addition, we observe that our method is very robust to $\lambda$ i.e.
-with subtle variations to differing $\lambda$, across the datasets.
-In particular, for CIFAR-10C, with $\lambda$ = 0 i.e.
-with no consistency loss, we achieve improvements of 0.37% and 3.86% over LAW [33] and TENT Continual [44].
-For CIFAR-100C, with no consistency loss ($\lambda$ = 0), we notice improvements of 1.31%, 2.06%, 4.96%, and 30.46% over LAW [33], CoTTA [45], BN Stats Adapt [37] and TENT Continual [44].
-As for ImageNet-C, the consistent best setting of $\lambda$ = 0.01 is better than LAW [33] by 1% and by 5.38% over Adacontrast [3].
-However, with $\lambda$ = 0 i.e.
-no consistency loss, we see a spike in the error rate by about 3.86% over the best setting ($\lambda$ = 0.01).
+Table 8 reports the average classification errors on the benchmark datasets.
+For $\lambda$ = 0.01, we achieve the best results.
+In addition, we observe that our method is very robust to $\lambda$ i.e., with subtle variations to differing $\lambda$.
 6 Conclusion
-In this paper, we propose a novel approach called PALM, to continual test-time adaptation.
-Our approach automatically decides on the importance of a layer by leveraging information from the gradient space i.e.
-by backpropagating the KL divergence between a uniform distribution and the softmax outputs.
+In this paper, we propose a novel approach called PALM, for continual test-time adaptation.
+Our approach automatically decides on the importance of a layer by leveraging information from the gradient space i.e., by backpropagating the KL divergence between a uniform distribution and the softmax outputs.
 This effectively eliminates the need for pseudo-labels.

### v3

@@
 Real-world vision models in dynamic environments face rapid shifts in domain distributions, leading to decreased recognition performance.
-Using unlabeled test data, continual test-time adaptation (CTTA) directly adjusts a pre-trained source discriminative model to these changing domains.
+Using unlabeled test data, continuous test-time adaptation (CTTA) directly adjusts a pre-trained source discriminative model to these changing domains.
 A highly effective CTTA method involves applying layer-wise adaptive learning rates for selectively adapting pre-trained layers.
@@
 We conduct extensive image classification experiments on CIFAR-10C, CIFAR-100C, and ImageNet-C, demonstrating the superior efficacy of our method compared to prior approaches.
+Codehttps://github.com/sarthaxxxxx/PALM
 1 Introduction
-In a real-world setting, machine-perception systems [1] function in a rapidly changing environment.
-In such contexts, pre-trained vision models are susceptible to performance degradation caused by potential distribution shifts.
+In a real-world setting, machine-perception systems (Arnold et al.
+2019) function in a rapidly changing environment.
+Pre-trained vision models are susceptible to performance degradation caused by potential distribution shifts in such contexts.
 For instance, while these models may have been trained on clean images, images captured by sensors could be corrupted by various weather conditions.
-In the realm of addressing domain shifts due to image corruptions, test-time adaptation (TTA) has been gaining a lot of traction lately [5, 36].
+In the realm of addressing domain shifts due to image corruptions, test-time adaptation (TTA) has been gaining a lot of traction lately (Döbler, Marsden, and Yang 2023; Wang et al.
+2020).
 Here, a pre-trained source model, with no access to its training (source) data due to privacy, is adapted to unlabeled test data through online training.
-One strategy in TTA involves mitigating domain shifts by adjusting the source model parameters using pseudo labels and minimizing the entropy of the model predictions [36].
+One strategy in TTA involves mitigating domain shifts by adjusting the source model parameters using pseudo labels and minimizing the entropy of the model predictions (Wang et al.
+2020).
 While this approach effectively adapts the model to a single type of corruption, it faces two constraints when continually applied across a sequence of test tasks featuring different corruptions.
@@
 Continual test-time adaptation (CTTA) has been recently introduced to tackle these challenges.
-In [23, 6, 36, 28], stable optimization is achieved by updating the batch normalization (BN) parameters or re-formulating its statistics.
-CoTTA [37] includes full model updates, with a teacher-student network setup, to handle the above issues.
-To further improve the performance of CoTTA, EcoTTA [33] introduces the attachment of “meta networks", i.e., a BN layer and a convolution block, to the frozen source model for efficient adaptation.
-More recently, the authors in [15] demonstrated that based on the type of distributional shift of the test data, a certain subset of layers of a pre-trained model can be “surgically" or heuristically chosen to fine-tune.
-Inspired by this finding, Layer-wise Auto-Weighting (LAW) [24] was introduced recently.
-LAW utilizes the Fisher Information Matrix (FIM) [27] to estimate domain shifts, automatically weighting trainable parameters and adjusting their learning rates (LRs) accordingly.
+In (Niu et al.
+2022; Gong et al.
+2022; Wang et al.
+2020; Schneider et al.
+2020), stable optimization is achieved by updating the batch normalization (BN) parameters or re-formulating its statistics.
+CoTTA (Wang et al.
+2022) includes full model updates, with a teacher-student network setup, to handle the above issues.
+To further improve the performance of CoTTA, EcoTTA (Song et al.
+2023) introduces the attachment of “meta networks”, i.e., a BN layer and a convolution block, to the frozen source model for efficient adaptation.
+More recently, the authors in (Lee et al.
+2022) demonstrated that based on the type of distributional shift of the test data, a certain subset of layers of a pre-trained model can be “surgically” or heuristically chosen to fine-tune.
+Inspired by this finding, Layer-wise Auto-Weighting (LAW) (Park et al.
+2024) was introduced recently.
+LAW utilizes the Fisher Information Matrix (FIM) (Sagun et al.
+2017) to estimate domain shifts, automatically weighting trainable parameters and adjusting their learning rates (LRs) accordingly.
 While LAW achieves state-of-the-art results for CTTA, it still encounters two significant limitations: 1) It presupposes that layer-wise importance can be accurately estimated via FIM with initially generated pseudo-labels.
@@
 2 Related Works
-Test-time adaptation (TTA): The goal of TTA is to adapt source models to unlabelled test data in real-time using a source-free online approach [12, 34, 36].
-TENT [36] optimizes BN parameters by minimizing the Shannon entropy [31] of pseudo labels.
-BN Stats Adapt (BN-1) [28] replaces the BN statistics with those from corrupted test data, while AdaContrast [3] uses contrastive learning and a memory module to update the entire model.
+Test-time adaptation (TTA): The goal of TTA is to adapt source models to unlabelled test data in real-time using a source-free online approach (Jain and Learned-Miller 2011; Sun et al.
+2020; Wang et al.
+2020).
+TENT (Wang et al.
+2020) optimizes BN parameters by minimizing the Shannon entropy (Shannon 1948) of pseudo labels.
+BN Stats Adapt (BN-1) (Schneider et al.
+2020) replaces the BN statistics with those from corrupted test data, while AdaContrast (Chen et al.
+2022) uses contrastive learning and a memory module to update the entire model.
 However, these methods address only single domain shifts and overlook continual shifts.
 Continual TTA (CTTA): CTTA requires models to adapt to dynamic domain shifts.
-CoTTA [37] uses a teacher-student setup with full model updates.
+CoTTA (Wang et al.
+2022) uses a teacher-student setup with full model updates.
 The teacher model is updated using an exponential moving average (EMA) to deal with catastrophic forgetting and error accumulation with time.
-EcoTTA [33] adds “meta networks" to a frozen source model, computing an $L_{1}$ regularization loss to prevent catastrophic forgetting.
-DSS [38] introduces a dynamic thresholding framework to differentiate between low and high sample quality.
-Efficient model fine-tuning: Freezing model parameters to learn information from downstream data has witnessed a good body of work [30, 17, 45, 7].
-Depending upon the distribution shift of the target task, in [15], the authors demonstrate that heuristically selecting a subset of convolutional blocks for training while freezing the remaining blocks works better than full fine-tuning.
-Adaptive LR methods: Research such as [40, 43] demonstrates that deeper layers of a network extract task-specific features, unlike shallow layers.
-AutoLR [26] developed an adaptive LR scheme based on weight variations between layers, although these rates tend to mostly stay constant during training.
-For CTTA, LAW [24] uses the FIM [27] to assess domain shifts and adjust parameter LRs accordingly.
+EcoTTA (Song et al.
+2023) adds “meta networks” to a frozen source model, computing an $L_{1}$ regularization loss to prevent catastrophic forgetting.
+DSS (Wang et al.
+2024) introduces a dynamic thresholding framework to differentiate between low and high sample quality.
+Efficient model fine-tuning: Freezing model parameters to learn information from downstream data has witnessed a good body of work (Sener et al.
+2016; Long et al.
+2016; Zintgraf et al.
+2019; Guo et al.
+2019).
+Depending upon the distribution shift of the target task, in (Lee et al.
+2022), the authors demonstrate that heuristically selecting a subset of convolutional blocks for training while freezing the remaining blocks works better than full fine-tuning.
+Adaptive LR methods: Research such as (Yosinski et al.
+2014; Zeiler and Fergus 2014) demonstrates that deeper layers of a network extract task-specific features, unlike shallow layers.
+AutoLR (Ro and Choi 2021) developed an adaptive LR scheme based on weight variations between layers, although these rates tend to mostly stay constant during training.
+For CTTA, LAW (Park et al.
+2024) uses the FIM (Sagun et al.
+2017) to assess domain shifts and adjust parameter LRs accordingly.
 However, using pseudo-labels to calculate FIM may lead to inaccuracies, and accumulating FIMs across domains can result in errors.
-Moreover, the adjustment of LR weights for each layer is constrained by the base LR, disregarding the relative “sensitivity" of parameters to the adaptation needs of the current batch and domain.
-Parameter sensitivity: Neural network pruning literature extensively explores parameter sensitivity [14, 18, 35, 20].
+Moreover, the adjustment of LR weights for each layer is constrained by the base LR, disregarding the relative “sensitivity” of parameters to the adaptation needs of the current batch and domain.
+Parameter sensitivity: Neural network pruning literature extensively explores parameter sensitivity (LeCun, Denker, and Solla 1989; Luo, Wu, and Lin 2017; Theis et al.
+2018; Molchanov et al.
+2019).
 It suggests that less sensitive parameters require higher LRs for effective adaptation, and vice-versa.
-Additionally, [16] introduces a concept of temporal sensitivity variation, influenced by uncertainty, as a factor in training models for tasks like natural language understanding and machine translation.
+Additionally, (Liang et al.
+2022) introduces a concept of temporal sensitivity variation, influenced by uncertainty, as a factor in training models for tasks like natural language understanding and machine translation.
 3 Preliminary
-3.1 Problem Setting
-CTTA aims to adapt a discriminative source model, parameterized as $\theta^{s}$ and trained on the source dataset $(\mathcal{X}^{\text{source}},\mathcal{Y}^{\text{source}})$, to multiple incoming tasks $\mathcal{T}_{i}=\{x_{k}\}_{k=1}^{K}$, at test-time.
+Problem setup: CTTA aims to adapt a discriminative source model, parameterized as $\theta^{s}$ and trained on the source dataset $(\mathcal{X}^{\text{source}},\mathcal{Y}^{\text{source}})$, to multiple incoming tasks $\mathcal{T}_{i}=\{x_{k}\}_{k=1}^{K}$, at test-time.
 Each $\mathcal{T}_{i}$ signifies the $i^{th}$ task with $K$ batches, across varying data distributions.
@@
 Each layer $\theta^{t}_{n}$ can represented by $\theta^{t}_{n}=\{\theta^{t}_{1,n},\ldots,\theta^{t}_{j,n},\ldots\}$, where $\theta_{j,n}^{t}$ is an arbitrary parameter of layer $\theta_{n}^{t}$.
-3.2 Core Motivations
-We provide a detailed view of surgical fine-tuning [15], AutoLR [26], and LAW [24].
+Core motivations: We provide a detailed view of surgical fine-tuning (Lee et al.
+2022), AutoLR (Ro and Choi 2021), and LAW (Park et al.
+2024).
 It is critical to understand the design principles and reasons for having an adaptive LR scheme for CTTA, which motivates this work.
@@
 Even though it is a desirable outcome for efficient fine-tuning, in this work, we close the gap by automatically selecting the layers that would adapt well to the target task.
-Inspired by [40, 43], AutoLR, proposed for supervised model fine-tuning on a downstream task, claims that the weight changes across layers should be in an increasing order, where the weight change $\delta w_{n}$ in the $n^{th}$ layer with weights $w_{n}$ at training epoch e is $\delta w_{n}=\|w_{n}^{e}\ -w_{n}^{(e-1)}\|$.
+Inspired by (Yosinski et al.
+2014; Zeiler and Fergus 2014), AutoLR, proposed for supervised model fine-tuning on a downstream task, claims that the weight changes across layers should be in an increasing order, where the weight change $\delta w_{n}$ in the $n^{th}$ layer with weights $w_{n}$ at training epoch e is $\delta w_{n}=\|w_{n}^{e}\ -w_{n}^{(e-1)}\|$.
 They hypothesize that initial layers, which extract general features, should have minimal $\delta w_{n}$, while deeper layers extract task-specific features.
@@
 However, the LRs for each layer show minimal deviation and remain consistent throughout the training process.
-As the first adaptive LR work for CTTA, LAW hypothesizes that identifying and controlling the adaptation rate of the different layers is useful [15].
+As the first adaptive LR work for CTTA, LAW hypothesizes that identifying and controlling the adaptation rate of the different layers is useful (Lee et al.
+2022).
 To understand the loss landscape and layer importance, the authors compute the gradient from the log-likelihood as $s(\theta_{j,n}^{t};x_{k})=\nabla_{\theta_{j,n}^{t}}\log(p(x_{k}))$, where $p(x_{k})$ is the predicted output.
-Then, the FIM [27], as a domain-shift indicator, is computed as a first-order approximation of the Hessian matrix.
+Then, the FIM (Sagun et al.
+2017), as a domain-shift indicator, is computed as a first-order approximation of the Hessian matrix.
 This layer-wise importance scheme completely relies on the pseudo-labels to compute the log-likelihood and is highly unreliable and uncertain.
 Due to the distribution shift, the label quality would begin to decrease for a continuum of tasks.
-The authors also propose to capture “domain-level" FIM ($\hat{F_{n}^{t}}$), for each layer, by a direct accumulation of the current layer-wise FIM (${F_{n}^{t}}$) as, $\hat{F}_{n}^{t}=\hat{F}_{n}^{t-1}+F_{n}^{t}$.
+The authors also propose to capture “domain-level” FIM ($\hat{F_{n}^{t}}$), for each layer, by a direct accumulation of the current layer-wise FIM (${F_{n}^{t}}$) as, $\hat{F}_{n}^{t}=\hat{F}_{n}^{t-1}+F_{n}^{t}$.
 This means that LAW puts equal emphasis on $\hat{F}_{n}^{t}$ and ${F_{n}^{t}}$, for each layer.
-An issue arises when ${F}_{n}^{t}$ of the current batch is small but gets outweighed by a “relatively" larger $\hat{F}_{n}^{t}$.
+An issue arises when ${F}_{n}^{t}$ of the current batch is small but gets outweighed by a “relatively” larger $\hat{F}_{n}^{t}$.
 This raises questions on the amount of adaptation required by the $n^{th}$ layer for batch $x_{k}$ as the FIMs dynamically change for each test batch.
@@
 We freeze the parameters of the unselected layers.
-4 Proposed Method
-Figure 1: PALM framework: At time t, input batch $x_{k}$ is processed by the model, parameterized by $\theta^{t}$.
-The KL-divergence between the softmax predictions and a uniform distribution is backpropagated to select the layers with the gradient norm $\leq$ $\eta$, to quantify the uncertainty.
-The parameter sensitivities of these layers, as an indicator of domain shift meters, are computed to update their learning rates (Eq.
-7).
-Finally, with the optimization objective in Eq.
-11, we update the model with the adjusted learning rates of the parameters.
-In this section, we introduce and formalize our proposed method for CTTA to address the key limitations of the works discussed above. Fig. 1 diagrammatically illustrates our entire approach.
-4.1 Layer Selection by Quantifying Prediction Uncertainty
-CTTA maximizes model performance in the current domain and so it is vital to assess prediction certainty amidst rapid domain shifts.
-In the literature on continual learning and TTA, several works utilize FIM [13, 29, 24].
-However, the unavailability of true labels, with an unknown distribution, pushes us to capture the model prediction uncertainty and adapt selected layers, automatically chosen, to a batch of test data, unlike surgical fine-tuning [15].
-Motivated by [11], which proposed an uncertainty scoring function from the gradient space, we compute the gradients of each layer.
+4 Proposed Methodology
+Figure 1: The framework of our proposed method PALM.
+Figure 1 illustrates our approach in two stages.
+Layer selection by quantifying prediction uncertainty: CTTA maximizes model performance in the current domain and so it is vital to assess prediction certainty amidst rapid domain shifts.
+In the literature on continual learning and TTA, several works utilize FIM (Kirkpatrick et al.
+2017; Schwarz et al.
+2018; Park et al.
+2024).
+However, the unavailability of true labels, with unknown distribution, pushes us to capture the model prediction uncertainty and adapt selected layers, automatically chosen, to a batch of test data, unlike surgical fine-tuning (Lee et al.
+2022).
+Motivated by (Huang, Geng, and Li 2021), which proposed an uncertainty scoring function from the gradient space, we compute the gradients of each layer.
 In essence, we backpropagate the KL divergence between the softmax probabilities and a uniform distribution u = [$\frac{1}{C}$, …, $\frac{1}{C}$] $\in$ $\mathbb{R}^{C}$, where $C$ is the total number of classes of the current task.
@@
 We freeze the layers with $\mathcal{Z}_{\theta_{j,n}^{t}}$ $>$ $\eta$ i.e., set their LR to 0.
-4.2 Parameter Sensitivity as a Domain Shift Indicator
-New studies have found that model parameters demonstrate levels of criticality [2, 44] across different positions.
-Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors [22, 25].
+Parameter sensitivity as a domain shift indicator: New studies have found that model parameters demonstrate levels of criticality (Chatterji, Neyshabur, and Sedghi 2019; Zhang, Bengio, and Singer 2022) across different positions.
+Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors (Mozer and Smolensky 1988; Rasmussen and Ghahramani 2000).
 Inspired by this observation, we assert that parameters contribute differently during CTTA.
@@
 For parameters of these selected layers only, we compute their sensitivity i.e., an approximation of an error when it is removed.
-Motivated by [14, 21, 20], we compute the sensitivity of a parameter $\theta_{j,n}^{t}$ based on the loss (Eq.
+Motivated by (LeCun, Denker, and Solla 1989; Molchanov et al.
+2016, 2019), we compute the sensitivity of a parameter $\theta_{j,n}^{t}$ based on the loss (Eq.
 1) as follows,
@@
 Precisely, $S_{j,n}^{t}$ approximates the change in the loss landscape at $\theta^{t}$ when $\theta_{j,n}^{t}$ is removed.
-Sensitivity, computed as a “weighting factor" of model weights [21, 20, 16], determines the level of adaptation needed for model parameters.
+Sensitivity, computed as a “weighting factor” of model weights (Molchanov et al.
+2016, 2019; Liang et al.
+2022), determines the level of adaptation needed for model parameters.
 We utilize the gradients of parameters from the selected layers, computed by Eq.
 2, to compactly derive Eq.
-3 from the first-order Taylor series expansion of $\mathcal{L}(\theta^{t})$ [lee2018snip] to,
+3 from the first-order Taylor series expansion of $\mathcal{L}(\theta^{t})$ (Lee, Ajanthan, and Torr 2018) to,
 $S_{j,n}^{t}=|\theta_{j,n}^{t}\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta^{t}})|$ (4)
@@
 However, $S_{j,n}^{t}$ of $\theta_{j,n}^{t}$ can be highly uncertain and would have a large variation.
-We introduce a method to capture this “domain-level" uncertainty by computing a weighted moving average of the sensitivity $S_{j,n}^{t}$.
+We introduce a method to capture this “domain-level” uncertainty by computing a weighted moving average of the sensitivity $S_{j,n}^{t}$.
 This is formulated as,
 $\hat{S}_{j,n}^{t}=\alpha S_{j,n}^{t}+(1-\alpha)\hat{S}_{j,n}^{(t-1)}$ (5)
-where $\hat{S}_{j,n}^{t}$ denotes the “domain-level" sensitivity, computed on-the-fly, and $\alpha$in [0,1] is a smoothing factor.
+where $\hat{S}_{j,n}^{t}$ denotes the “domain-level” sensitivity, computed on-the-fly, and $\alpha$in [0,1] is a smoothing factor.
 Specifically, $\alpha$ decides the contribution of the current sensitivity $S_{j,n}^{t}$ to $\hat{S}_{j,n}^{t}$.
 Since CTTA happens at the batch level, it is critical to achieve a good balance between $S_{j,n}^{t}$ and $\hat{S}_{j,n}^{t}$, with a slightly larger focus on $S_{j,n}^{t}$.
-As can be seen, we do not directly accumulate the “domain-level" information, unlike LAW.
+As can be seen, we do not directly accumulate the “domain-level” information, unlike LAW.
 Due to large distribution shifts happening because of the continuum of tasks, larger deviations of $S_{n}^{t}$ from $\hat{S}_{n}^{t}$ are bound to happen.
@@
 This would take care of error accumulation too.
-To automatically alter the LRs of these parameters, the “importance" ($i_{j,n}^{t}$) is calculated as,
+To automatically alter the LRs of these parameters, the “importance” ($i_{j,n}^{t}$) is calculated as,
 $i_{j,n}^{t}=\hat{D}_{j,n}^{t}/\hat{S}_{j,n}^{t}$ (7)
@@
 7.
-4.3 Optimization Objective
-Though our method is robust, it is necessary to avoid miscalibrated labels.
-TENT [36] minimized the entropy on the output logits, l, following $\mathcal{H}(l)=-\sum_{c}p(l_{c})\log(p(l_{c}))$, where $p(l_{c})$ is the probability for class c.
-Following EcoTTA [33], we use entropy minimization as the objective for adaptation where $\mathcal{H}(l)$ is minimized for samples below a certain fixed threshold $\mathcal{H}_{0}$ as,
+Optimization: Though our method is robust, it is necessary to avoid miscalibrated labels.
+TENT (Wang et al.
+2020) minimized the entropy on the output logits, l, following $\mathcal{H}(l)=-\sum_{c}p(l_{c})\log(p(l_{c}))$, where $p(l_{c})$ is the probability for class c.
+Following EcoTTA (Song et al.
+2023), we use entropy minimization as the objective for adaptation where $\mathcal{H}(l)$ is minimized for samples below a certain fixed threshold $\mathcal{H}_{0}$ as,
 $\mathcal{L}^{entropy}=\mathbf{1}_{\mathcal{H}(l)\leq\mathcal{H}_{0}}\mathcal{H}(l)$ (9)
 $\mathcal{H}_{0}$ is set to $0.4\times\log C$ following EcoTTA.
-For a fair comparison with prior works [37, 24], we apply a consistency loss between l and $\hat{l}$, computed as the output logits from the augmented version of $x_{k}$, to regularize the model as follows,
+For a fair comparison with prior works (Wang et al.
+2022; Park et al.
+2024), we apply a consistency loss between l and $\hat{l}$, computed as the output logits from the augmented version of $x_{k}$, to regularize the model as follows,
 $\mathcal{L}^{const}=-\sum_{c}p(l_{c})\log(p(\hat{l}_{c}))$ (10)
@@
 where $\lambda$ is a regularisation coefficient. In summary, upon calculating the parameter-specific LRs (Eq. 8), we optimize the continually fine-tuned model to minimize Eq. 11.
-5 Experiments
-We draw a contrast of our proposed method against other state-of-the-art works in a continual test-time adaptation (CTTA) setting.
-However, to have a broader study and analysis, we also report results in a gradual test-time adaptation (GTTA) setting, following [37, 24].
-In the following subsections, we describe the implementation details of both the challenging adaptation settings.
-5.1 Experimental Settings and Baselines
-Datasets and tasks.
-Following the standard benchmarks set by CoTTA [37], we evaluate our proposed method on CIFAR-10C, CIFAR-100C, and ImageNet-C, based on image corruption schemes as set in [9].
-Each dataset contains a set of 15 corruption styles as tasks (e.g.
-gaussian noise, shot noise, impulse noise, $\ldots$) with 5 severity levels.
-We follow CoTTA and apply the same test-time augmentations, like color jitter, gaussian blur, gaussian noise, random flips, and random affine, to the input $x_{k}$.
-As the evaluation metric, we report the mean classification error.
-CTTA implementation details.
-We evaluate our approach following prior state-of-the-art methods [36, 37, 3, 33, 28, 24].
-To maintain fairness regarding the usage of pre-trained source models, we employ WideResNet-28 [42] for CIFAR-10C, ResNeXt-29 [39] for CIFAR-100C, and ResNet-50 [8] for ImageNet-C, all available on RobustBench [4].
-We optimize using an Adam optimizer, setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C.
-For balancing parameter sensitivities, we set $\alpha$ to 0.5, 0.9, and 0.5 respectively, with temperature coefficients T set to 50, 100, and 1000 respectively.
-We set $\eta$ to 1, 0.5, and 0.3, and $\lambda$ to 0.01 throughout.
-Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA, and results are only reported for the highest severity level of 5 for each task.
-GTTA implementation details.
-In the standard CTTA setup, each task only receives images with the highest severity.
-However, in GTTA, following CoTTA, for each task, the severity levels are gradually changed as $\ldots$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5$\rightarrow$4$\rightarrow$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\ldots$.
-We adhere to the details specified in the CTTA setting.
-Baselines.
-We compare against several CTTA baselines, including BN Stats Adapt (BN-1) [28], TENT Continual [36], CoTTA [37], EcoTTA [33], AdaContrast [3], DSS [38], and LAW [24].
-We also contrast our work against surgical fine-tuning [15].
-For EcoTTA, we report the results on CIFAR-10C and ImageNet-C only since they do not use ResNeXt-29 [39] for CIFAR-100C.
-Following the findings of surgical fine-tuning [15], we continually fine-tune only the first convolutional block of the respective models, denoted as “Surgical fine-tuning (continual)", using an SGD optimizer with fixed learning rates of 1e-3 for the CIFAR datasets and 5e-4 for ImageNet-C.
-For the reported results of LAW, we reproduce their results strictly following their official implementation [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
+5 Experiments and Results
 Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
@@
 Source 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.5 30.3 43.5
-BN-1 [28] 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
-TENT Continual [36] 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
-CoTTA [37]
+BN-1 (Schneider et al. 2020) 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
+TENT Continual (Wang et al. 2020) 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
+CoTTA (Wang et al. 2022)
 24.3 21.3 26.6 11.6 27.6 12.2 10.3 14.8 14.1 12.4 7.5 10.6 18.3 13.4 17.3 16.2
-EcoTTA [33]
+EcoTTA (Song et al. 2023)
 23.8 18.7 25.7 11.5 29.8 13.3 11.3 15.3 15.0 13.0 7.9 11.3 20.2 15.1 20.5 16.8
-AdaContrast [3]
+AdaContrast (Chen et al. 2022)
 29.1 22.5 30.0 14.0 32.7 14.1 12.0 16.6 14.9 14.4 8.1 10.0 21.9 17.7 20.0 18.5
 Surgical
-fine-tuning [15]
+fine-tuning (Lee et al. 2022)
 27.7 24.1 33.8 12.2 33.1 13.1 10.9 15.9 15.6 13.8 7.8 10.8 21.2 16.6 22.9 18.6
-DSS [38]
+DSS (Wang et al. 2024)
 24.1 21.3 25.4 11.7 26.9 12.2 10.5 14.5 14.1 12.5 7.8 10.8 18.0 13.1 17.3 16.0
-LAW [24]
+LAW (Park et al. 2024)
 24.7 18.9 25.5 12.9 26.7 15.0 11.8 15.1 14.7 15.9 10.1 13.8 19.4 14.7 18.3 17.2
@@
 73.0 68 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
-BN-1 [28]
+BN-1 (Schneider et al. 2020)
 42.1 40.7 42.7 27.6 41.9 29.7 27.9 34.9 35.0 41.5 26.5 30.3 35.7 32.9 41.2 35.4
-TENT Continual [36]
+TENT Continual (Wang et al. 2020)
 37.2 35.8 41.7 37.9 51.2 48.3 48.5 58.4 63.7 71.1 70.4 82.3 88.0 88.5 90.4 60.9
-CoTTA [37]
+CoTTA (Wang et al. 2022)
 40.1 37.7 39.7 26.9 38.0 27.9 26.4 32.8 31.8 40.3 24.7 26.9 32.5 28.3 33.5 32.5
-AdaContrast [3]
+AdaContrast (Chen et al. 2022)
 42.3 36.8 38.6 27.7 40.1 29.1 27.5 32.9 30.7 38.2 25.9 28.3 33.9 33.3 36.2 33.4
 Surgical
-fine-tuning [15]
+fine-tuning (Lee et al. 2022)
 40.7 36.3 37.8 25.5 39.0 27.6 25.2 32.2 30.4 37.3 23.7 26.8 32.3 27.6 37.2 32.0
-DSS[38]
+DSS(Wang et al. 2024)
 39.7 36.0 37.2 26.3 35.6 27.5 25.1 31.4 30.0 37.8 24.2 26.0 30.0 26.3 31.1 30.9
-LAW [24]
+LAW (Park et al. 2024)
 41.0 36.7 38.3 25.6 37.0 27.8 25.2 30.7 30.0 37.3 24.4 27.6 31.2 27.8 34.9 31.7
@@
 Table 2: Mean errors (%) on CIFAR-100C - CTTA mean errors of the 15 corruptions (tasks) at a severity level of 5.
-5.2 CTTA Classification Results
-Results on CIFAR-10C and CIFAR-100C.
-We contrast our method against all the baselines as in Tables 1 and 2.
-The results reveal that our method performs the best across all training setups and baselines for both datasets.
-We achieve large improvements of 4.9% and 5.2% for CIFAR-10C and 5.3% and 30.8% for CIFAR-100C, over BN-1 and TENT Continual respectively.
-Concerning recent works of LAW and DSS, our method improves 1.7% and 0.5% on CIFAR-10C and 1.6% and 0.8% on CIFAR-100C, respectively.
-It is noteworthy that TENT Continual initially excels in the early tasks but experiences error accumulation with time, leading to an increase in mean error.
-In the study by CoTTA, the method’s reliance on full model updates renders it computationally expensive and unsuitable for deployment.
-An interesting observation is that with a drastic reduction in the number of trainable parameters, as in “Surgical fine-tuning (continual)", the mean errors are better than most CTTA benchmarks, if not the best.
-LAW uses the uncertain pseudo labels to calculate the domain shift approximated via the FIM.
-The minor error that is introduced for each batch begins to propagate across all the domains.
-In addition, the imbalance caused by the “domain-level" FIM and the current batch’s FIM, results in a higher error rate for the initial tasks.
-In our method, the computation of the model prediction uncertainty to automatically select layers is useful (Eq.
-1).
-This affirms the findings in [15] that concentrated adaptation, for image corruptions, in the initial layers suffices.
-Results on Imagenet-C In Table 3, we report the CTTA results on the ImageNet-C dataset.
-As shown, our method outperforms all the previous CTTA baselines and “Surgical fine-tuning (continual)".
-We observe an improvement of 0.9% and 4.5% over LAW and DSS respectively.
-In addition, we achieve 2.6%, 5.4%, and 2.5% improvements over CoTTA, AdaContrast, and TENT Continual, respectively.
+We evaluate our proposed method in a CTTA setting, comparing it to state-of-the-art approaches, and extend our study to the gradual test-time adaptation (GTTA) setting following (Wang et al.
+2022; Park et al.
+2024).
+Experimental Settings and Baselines
+Datasets and Tasks: Following the standard benchmarks set by (Wang et al.
+2022), we evaluate our proposed method on CIFAR-10C, CIFAR-100C, and ImageNet-C, based on image corruption schemes as set in (Hendrycks and Dietterich 2019).
+Each dataset contains a set of 15 corruption styles as tasks (e.g.
+gaussian noise, shot noise, $\ldots$) with 5 severity levels.
+We follow CoTTA and apply the same test-time augmentations, like color jitter, gaussian blur, gaussian noise, random flips, and random affine, to the input $x_{k}$.
+As the evaluation metric, we report the mean classification error.
+CTTA implementation details: To maintain fairness regarding the source models, we employ WideResNet-28 (Zagoruyko and Komodakis 2016) for CIFAR-10C, ResNeXt-29 (Xie et al.
+2017) for CIFAR-100C, and ResNet-50 (He et al.
+2016) for ImageNet-C, all available on RobustBench (Croce et al.
+2020).
+We optimize using an Adam optimizer, setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C.
+For balancing parameter sensitivities, we set $\alpha$ to 0.5, 0.9, and 0.5 respectively, with temperature coefficients T set to 50, 100, and 1000 respectively.
+We set $\eta$ to 1, 0.5, and 0.3, and $\lambda$ to 0.01 throughout.
+Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA, and results are only reported for the highest severity level of 5 for each task.
+GTTA implementation details: In the standard CTTA setup, each task only receives images with the highest severity.
+However, in GTTA, following CoTTA, for each task, the severity levels are gradually changed as $\ldots$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5$\rightarrow$4$\rightarrow$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\ldots$.
+We adhere to the details specified in the CTTA setting.
+Baselines: We compare against several CTTA baselines, including BN Stats Adapt (BN-1) (Schneider et al.
+2020), TENT Continual (Wang et al.
+2020), CoTTA (Wang et al.
+2022), EcoTTA (Song et al.
+2023), AdaContrast (Chen et al.
+2022), DSS (Wang et al.
+2024), and LAW (Park et al.
+2024).
+We also contrast our work against surgical fine-tuning (Lee et al.
+2022).
+For EcoTTA, we report the results on CIFAR-10C and ImageNet-C only since they do not use ResNeXt-29 (Xie et al.
+2017) for CIFAR-100C.
+Following the findings of (Lee et al.
+2022), we continually fine-tune only the first convolutional block of the respective models, denoted as “Surgical fine-tuning (continual)”, using an SGD optimizer with fixed learning rates of 1e-3 for the CIFAR datasets and 5e-4 for ImageNet-C.
+For the results of LAW, we reproduce their results strictly following their official implementation  [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
 Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
@@
 97.8 97.1 98.2 81.7 89.8 85.2 78.0 83.5 77.1 75.9 41.3 94.5 82.5 79.3 68.6 82.0
-BN-1 [28]
+BN-1 (Schneider et al. 2020)
 85.0 83.7 85.0 84.7 84.3 73.7 61.2 66.0 68.2 52.1 34.9 82.7 55.9 51.3 59.8 68.6
-TENT Continual [36]
+TENT Continual (Wang et al. 2020)
 81.6 74.6 72.7 77.6 73.8 65.5 55.3 61.6 63.0 51.7 38.2 72.1 50.8 47.4 53.3 62.6
-CoTTA [37]
+CoTTA (Wang et al. 2022)
 84.7 82.1 80.6 81.3 79.0 68.6 57.5 60.3 60.5 48.3 36.6 66.1 47.2 41.2 46.0 62.7
-EcoTTA [33]
+EcoTTA (Song et al. 2023)
 - - - - - - - - - - - - - - - 63.4
-AdaContrast [3]
+AdaContrast (Chen et al. 2022)
 82.9 80.9 78.4 81.4 78.7 72.9 64.0 63.5 64.5 53.5 38.4 66.7 54.6 49.4 53.0 65.5
 Surgical
-fine-tuning [15]
+fine-tuning (Lee et al. 2022)
 83.4 79.5 77.2 86.1 82.4 72.2 61.1 63.1 64.7 51.2 35.9 72.9 56.2 46.7 50.2 65.5
-DSS [38]
+DSS (Wang et al. 2024)
 84.6 80.4 78.7 83.9 79.8 74.9 62.9 62.8 62.9 49.7 37.4 71.0 49.5 42.9 48.2 64.6
-LAW [24]
+LAW (Park et al. 2024)
 81.8 75.0 72.1 77.4 73.2 63.9 54.6 57.9 61.0 50.1 36.3 68.6 49.0 46.2 49.1 61.0
@@
 Table 3: Mean errors (%) on ImageNet-C - CTTA mean errors of the 15 corruptions (tasks) at a severity level of 5.
+CTTA Classification Results
+Results on CIFAR-10C and CIFAR-100C: We report the CTTA results in Tables 1 and 2.
+Our method performs the best across all training setups and baselines for both datasets.
+We achieve large improvements of 4.9% and 5.2% for CIFAR-10C and 5.3% and 30.8% for CIFAR-100C, over BN-1 and TENT Continual respectively.
+Concerning recent works of LAW and DSS, our method improves by 1.7% and 0.5% on CIFAR-10C and 1.6% and 0.8% on CIFAR-100C, respectively.
+TENT Continual excels initially but accumulates errors over time, increasing mean error.
+In CoTTA, the method’s reliance on full model updates renders it computationally expensive and unsuitable for deployment.
+An interesting observation is that with a drastic reduction in the number of trainable parameters, as in “Surgical fine-tuning (continual)”, the mean errors are better than most CTTA benchmarks, if not the best.
+LAW uses the uncertain pseudo labels to calculate the domain shift approximated via the FIM.
+The minor error that is introduced for each batch begins to propagate across all the domains.
+In addition, the imbalance caused by the “domain-level” FIM and the current batch’s FIM, results in a higher error rate for the initial tasks.
+In our method, the computation of the model prediction uncertainty to automatically select layers is useful (Eq.
+1).
+This affirms the findings in (Lee et al.
+2022) that concentrated adaptation, for image corruptions, in the initial layers suffices.
+Results on Imagenet-C: In Table 3, we report the CTTA results on the ImageNet-C dataset.
+As shown, our method outperforms all the previous CTTA baselines and “Surgical fine-tuning (continual)”.
+We observe an improvement of 0.9% and 4.5% over LAW and DSS respectively.
+In addition, we achieve 2.6%, 5.4%, and 2.5% improvements over CoTTA, AdaContrast, and TENT Continual, respectively.
 Dataset Source
@@
 Table 4: Mean errors (in %) on CIFAR-10C, CIFAR-100C, and ImageNet-C - We report the GTTA mean errors averaged across all the severity levels (1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5).
-5.3 GTTA Classification Results
-Results on CIFAR-10C, CIFAR-100C, and Imagenet-C.
-Here, we present the gradual test-time adaptation (GTTA) results on CIFAR-10C, CIFAR-100C, and Imagenet-C.
+GTTA Classification Results
+Results: Here, we present the GTTA results on the benchmark datasets.
 For each task, the corruption severity of the inputs is gradually changed as discussed earlier.
@@
 With most parameters frozen during the adaptation on each batch, our method is more robust and inexpensive than CoTTA, which also demonstrates competitive improvements in the mean classification error.
-5.4 Ablation Studies
-We conduct ablation experiments to validate our approach and gain insights into its underlying mechanisms and components.
-Specifically, we conduct these experiments and report results on the three CTTA benchmark datasets: CIFAR-10C, CIFAR-100C, and ImageNet-100C.
-    WideResNet-28 [42]       ResNeXt-29 [39]       ResNet-50 [8]
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Figure 2: Illustrations of LR importance $i_{j,n}^{t}$ across different convolutional blocks/stages/layers during CTTA.
-[Top row] - Variations for the domain “glass_blur" across datasets.
-[Bottom row] - Variations for the domain “snow".
+Ablation Studies and Analysis
+Computational efficiency and Catastrophic forgetting (CF): In our PALM, we first select a subset of layers from the gradient space (Eq.
+2) and adapt them while strictly freezing the unselected layers.
+In Table 5, we report the average % of adapted model parameters per task with respect to the source model, for our method and CoTTA, across the benchmark datasets.
+Our method is computationally inexpensive and on an average updates 4.34%, 6.65%, and 17.21% of their respective source parameters.
+This showcases significant computational advantages over CoTTA, which updates 100% of the parameters.
+With a single NVIDIA A5000 GPU, PALM incurs a slightly higher adaptation time/batch, due to the two-stage proposed method.
+For example, on CIFAR-10, PALM runs for 0.8s per batch, whereas Surgical Fine-tuning (Surgical) and LAW take 0.45 s and 0.73 s respectively.
+Our method alleviates CF by only adapting a small subset of model parameters, ensuring that much of the prior source knowledge is preserved.
+Quantitatively, the source models achieved zero-shot mean accuracies of 94.03%, 76.06%, and 75.19% on respective datasets.
+After adaptation, PALM retained source knowledge better, with mean accuracies of 92.29%, 77.46%, and 70.87%, compared to LAW’s 91.67%, 77.69%, and 71.57%.
+Despite accuracy decline with more tasks, especially after adaptation on ImageNet-C, PALM matches or outperforms LAW.
 Dataset Method
@@
 Table 5: Average % of adapted parameters/task.
-Computational efficiency and Catastrophic forgetting: In our proposed method, we first select a subset of layers from the gradient space (Eq.
-2) and adapt them while strictly freezing the unselected layers.
-In Table 5, we report the average % of adapted model parameters per task with respect to the source model, for our method and CoTTA, across the benchmark datasets.
-Our method is computationally inexpensive and on an average updates 4.34%, 6.65%, and 17.21% of their respective source model parameters.
-This showcases significant computational advantages over CoTTA, which updates 100% of the source parameters.
-Additionally, in LAW, the LRs of all the parameters are adjusted.
-It is worth mentioning that our method alleviates catastrophic forgetting (CF) by only adapting a small subset of model parameters ensuring that much of the prior source knowledge is preserved and CF is largely minimized.
+Adaptive LR optimizers vs Ours: Here, we study the effect of directly using adaptive LR optimizers like LARS (You, Gitman, and Ginsburg 2017), Prodigy (Mishchenko and Defazio 2023), and Adafactor (Shazeer and Stern 2018), for CTTA.
+We report the results in Table 6.
+The mentioned optimizers do not adjust layer-wise LRs based on parameter sensitivity, failing to capture rapid domain shifts, resulting in poorer performance compared to our method.
 Dataset Ours LARS Prodigy Adafactor
@@
 Table 6: CTTA mean errors (%) using adaptive LR optimizers vs. ours.
-Adaptive LR optimizers vs Ours: Here, we study the effect of directly using adaptive LR optimizers like LARS [41], Prodigy [19], and Adafactor [32], for CTTA.
-We report the results in Table 6.
-The mentioned optimizers do not adjust layer-wise LRs based on parameter sensitivity, failing to capture rapid domain shifts, resulting in poorer performance compared to our method.
+    WideResNet-28     ResNeXt-29     ResNet-50
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Figure 2: Illustrations of LR importance $i_{j,n}^{t}$ across different convolutional blocks/stages/layers during CTTA.
+[Top row] - Variations for the domain “glass_blur” across datasets.
+[Bottom row] - Variations for the domain “snow”.
 Analysis of LR importance $i_{j,n}^{t}$: The prime objective of our work is to adaptively alter the LRs of parameters of selected layers based on their sensitivity.
@@
 Across all the models, we notice that generally, the deeper layers have smaller associated $i_{j,n}^{t}$, which means that the initial layers are adapted more to the target domain.
-This result completely aligns with the findings of surgical fine-tuning [15].
+This result completely aligns with the findings of surgical fine-tuning (Lee et al.
+2022).
 In addition to that, we also notice more sparsity in the deeper layers.
 This further confirms the necessity of freezing parameters automatically selected based on the target domain for adaptation.
-Layer importance via pseudo-labels: In LAW, the authors use the pseudo-labels to decide on the importance of layers for adaptation.
-In this, we study the effectiveness of such a design for our method i.e., we investigate if layer importance computed using pseudo-labels is helpful.
+Layer importance via pseudo-labels: LAW leverages pseudo-labels to determine layer importance for adaptation.
+We evaluate the effectiveness of this design for our method by investigating the utility of pseudo-label-based layer importance.
 We compute the gradients from the log-likelihood as $\nabla_{\theta_{j,n}^{t}}\log(p(x_{k}))$ and estimate the sensitivity of the parameters of the chosen layers (Eq.
 4-Eq.
-6), and alter the learning rates (Eq.
+6), and alter the LRs (Eq.
 7).
 Table 7 reports the mean classification errors for all the datasets.
-We observe that the gradients computed using the pseudo-labels are an unreliable indicator of prediction uncertainty.
+We observe that such gradients are unreliable indicators of prediction uncertainty.
 The scoring function $\mathcal{Z}_{\theta_{j,n}^{t}}=\|{{\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta}^{t})}}\|_{1}$ computed using these gradients carries insufficient information.
@@
 Refer to caption Refer to caption
-a) Ablation on smoothing factor $\alpha$. b) Ablation on temperature T.
-Figure 3: Ablation results on smoothing factor $\alpha$ and temperature T for CIFAR-10C, CIFAR-100C, and ImageNet-C.
+a) Ablation on $\alpha$. b) Ablation on temperature T.
+Figure 3: Ablation results on smoothing factor $\alpha$ and temperature T.
 Dataset Pseudo-labels Ours
@@
 Table 8: Ablation results with varying $\lambda$.
-Contribution of smoothing factor $\alpha$: A key component of our method is computing the moving average of sensitivities across data batches, smoothed by parameter $\alpha$ (Eq.
+Contribution of smoothing factor $\alpha$: A key component of PALM is computing the moving average of sensitivities across data batches, smoothed by $\alpha$ (Eq.
 5).
-Fig.
-3(a) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$.
-We observe that the “domain-level" sensitivity across past batches of data is indeed useful.
+Figure 3(a) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$.
+We observe that the “domain-level” sensitivity across past batches of data is indeed useful.
 For $\alpha$ = 0 i.e., when $S_{j,n}^{t}$ of a selected parameter at timestep t is not considered, the respective datasets’ error rate drastically spikes.
-This suggests that the effect of the current sensitivity of selected layers is indeed required.
 For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.7% over LAW.
@@
 Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{j,n}^{t}$) and current sensitivity ($S_{j,n}^{t}$) is needed.
-[…]
+Effectiveness of temperature T: In our method, we compute the KL divergence between a uniform distribution u and a smoothed classifier’s output logits (Eq.
+1), denoted by $\hat{h}(x_{k})$.
+To study its influence, we illustrate the results in Figure 3(b).
+Ideally, larger values of T push the distributions of the predictions closer to u (Hinton, Vinyals, and Dean 2015) by spreading the probability among inputs.
+In CTTA, since we deal with domain shifts with no labels, we argue that higher values of T are useful to capture the uncertainty.
+So, the KL divergence reduces, indicating more uncertainty in the model and hence, more adaptation.
+For a systematic temperature selection, we aim to adjust T based on the number of adapted parameters.
+In particular, if the number of adapted parameters increases across batches, T will be increased.
+For a difficult dataset like ImageNet-C, larger values of T are useful with the error reducing as T increases.
+For the CIFAR-C datasets, the errors begin to rise for T $\geq$ 100 which could be due to large spreads of the distributions.
 Robustness to regularization $\lambda$: Here, we understand the effect of the parameter $\lambda$ in Eq.
@@
 The results reveal that PALM achieves superior performance and pushes the state-of-the-art on several standard baselines.
+Acknowledgments. We thank all the reviewers for their helpful comments and suggestions. This project was supported by a grant from the University of Texas at Dallas, Richardson, TX, USA.

### v4

@@
 Codehttps://github.com/sarthaxxxxx/PALM
-1 Introduction
+Introduction
 In a real-world setting, machine-perception systems (Arnold et al.
@@
 3) Our method delivers superior results, via extensive evaluations, in both continual and gradual TTA across three standard benchmark datasets.
-2 Related Works
+Related Works
 Test-time adaptation (TTA): The goal of TTA is to adapt source models to unlabelled test data in real-time using a source-free online approach (Jain and Learned-Miller 2011; Sun et al.
@@
 2022) introduces a concept of temporal sensitivity variation, influenced by uncertainty, as a factor in training models for tasks like natural language understanding and machine translation.
-3 Preliminary
+Preliminary
 Problem setup: CTTA aims to adapt a discriminative source model, parameterized as $\theta^{s}$ and trained on the source dataset $(\mathcal{X}^{\text{source}},\mathcal{Y}^{\text{source}})$, to multiple incoming tasks $\mathcal{T}_{i}=\{x_{k}\}_{k=1}^{K}$, at test-time.
@@
 Each layer $\theta^{t}_{n}$ can represented by $\theta^{t}_{n}=\{\theta^{t}_{1,n},\ldots,\theta^{t}_{j,n},\ldots\}$, where $\theta_{j,n}^{t}$ is an arbitrary parameter of layer $\theta_{n}^{t}$.
-Core motivations: We provide a detailed view of surgical fine-tuning (Lee et al.
+Core motivations: We provide a detailed view of surgical fine-tuning (SFT) (Lee et al.
 2022), AutoLR (Ro and Choi 2021), and LAW (Park et al.
@@
 It is critical to understand the design principles and reasons for having an adaptive LR scheme for CTTA, which motivates this work.
-To preserve the information during the source pre-training phase while fine-tuning on a target task of a different distribution like an image corruption, surgical fine-tuning empirically shows that heuristically fine-tuning only the first convolutional block results in better performance over full fine-tuning.
+To preserve the information during the source pre-training phase while fine-tuning on a target task of a different distribution like an image corruption, SFT empirically shows that heuristically fine-tuning only the first convolutional block results in better performance over full fine-tuning.
 Even though it is a desirable outcome for efficient fine-tuning, in this work, we close the gap by automatically selecting the layers that would adapt well to the target task.
@@
 We freeze the parameters of the unselected layers.
-4 Proposed Methodology
+Proposed Methodology
 Figure 1: The framework of our proposed method PALM.
@@
 2024).
-However, the unavailability of true labels, with unknown distribution, pushes us to capture the model prediction uncertainty and adapt selected layers, automatically chosen, to a batch of test data, unlike surgical fine-tuning (Lee et al.
+However, the unavailability of true labels, with unknown distribution, pushes us to capture the model prediction uncertainty and adapt selected layers, automatically chosen, to a batch of test data, unlike SFT (Lee et al.
 2022).
@@
 We freeze the layers with $\mathcal{Z}_{\theta_{j,n}^{t}}$ $>$ $\eta$ i.e., set their LR to 0.
-Parameter sensitivity as a domain shift indicator: New studies have found that model parameters demonstrate levels of criticality (Chatterji, Neyshabur, and Sedghi 2019; Zhang, Bengio, and Singer 2022) across different positions.
+Parameter sensitivity as a domain shift indicator: New studies have found that model parameters demonstrate levels of criticality (Zhang, Bengio, and Singer 2022) across different positions.
 Moreover, not all parameters adapt uniformly, and removing some can lower test generalization errors (Mozer and Smolensky 1988; Rasmussen and Ghahramani 2000).
@@
 For parameters of these selected layers only, we compute their sensitivity i.e., an approximation of an error when it is removed.
-Motivated by (LeCun, Denker, and Solla 1989; Molchanov et al.
+Motivated by (Molchanov et al.
 2016, 2019), we compute the sensitivity of a parameter $\theta_{j,n}^{t}$ based on the loss (Eq.
@@
 where $\lambda$ is a regularisation coefficient. In summary, upon calculating the parameter-specific LRs (Eq. 8), we optimize the continually fine-tuned model to minimize Eq. 11.
-5 Experiments and Results
+Experiments and Results
 Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Brightness Contrast
@@
 Source 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.5 30.3 43.5
-BN-1 (Schneider et al. 2020) 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
-TENT Continual (Wang et al. 2020) 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
-CoTTA (Wang et al. 2022)
-24.3 21.3 26.6 11.6 27.6 12.2 10.3 14.8 14.1 12.4 7.5 10.6 18.3 13.4 17.3 16.2
-EcoTTA (Song et al. 2023)
-23.8 18.7 25.7 11.5 29.8 13.3 11.3 15.3 15.0 13.0 7.9 11.3 20.2 15.1 20.5 16.8
-AdaContrast (Chen et al. 2022)
-29.1 22.5 30.0 14.0 32.7 14.1 12.0 16.6 14.9 14.4 8.1 10.0 21.9 17.7 20.0 18.5
-Surgical
-fine-tuning (Lee et al. 2022)
-27.7 24.1 33.8 12.2 33.1 13.1 10.9 15.9 15.6 13.8 7.8 10.8 21.2 16.6 22.9 18.6
-DSS (Wang et al. 2024)
-24.1 21.3 25.4 11.7 26.9 12.2 10.5 14.5 14.1 12.5 7.8 10.8 18.0 13.1 17.3 16.0
-LAW (Park et al. 2024)
-24.7 18.9 25.5 12.9 26.7 15.0 11.8 15.1 14.7 15.9 10.1 13.8 19.4 14.7 18.3 17.2
-Ours
-25.8 18.1 22.7 12.3 25.3 13.1 10.7 13.5 13.1 12.2 8.5 11.8 17.9 12.0 15.4 15.5
+BN-1 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
+Tent Cont. 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
+CoTTA 24.3 21.3 26.6 11.6 27.6 12.2 10.3 14.8 14.1 12.4 7.5 10.6 18.3 13.4 17.3 16.2
+EcoTTA 23.8 18.7 25.7 11.5 29.8 13.3 11.3 15.3 15.0 13.0 7.9 11.3 20.2 15.1 20.5 16.8
+AdaContrast 29.1 22.5 30.0 14.0 32.7 14.1 12.0 16.6 14.9 14.4 8.1 10.0 21.9 17.7 20.0 18.5
+SFT (cont.) 27.7 24.1 33.8 12.2 33.1 13.1 10.9 15.9 15.6 13.8 7.8 10.8 21.2 16.6 22.9 18.6
+DSS 24.1 21.3 25.4 11.7 26.9 12.2 10.5 14.5 14.1 12.5 7.8 10.8 18.0 13.1 17.3 16.0
+LAW 24.7 18.9 25.5 12.9 26.7 15.0 11.8 15.1 14.7 15.9 10.1 13.8 19.4 14.7 18.3 17.2
+Ours 25.8 18.1 22.7 12.3 25.3 13.1 10.7 13.5 13.1 12.2 8.5 11.8 17.9 12.0 15.4 15.5
 Table 1: Mean errors (%) on CIFAR-10C - CTTA mean errors of the 15 corruptions (tasks) at a severity level of 5.
@@
 Pixelate JPEG Mean
-Source
-73.0 68 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
-BN-1 (Schneider et al. 2020)
-42.1 40.7 42.7 27.6 41.9 29.7 27.9 34.9 35.0 41.5 26.5 30.3 35.7 32.9 41.2 35.4
-TENT Continual (Wang et al. 2020)
-37.2 35.8 41.7 37.9 51.2 48.3 48.5 58.4 63.7 71.1 70.4 82.3 88.0 88.5 90.4 60.9
-CoTTA (Wang et al. 2022)
-40.1 37.7 39.7 26.9 38.0 27.9 26.4 32.8 31.8 40.3 24.7 26.9 32.5 28.3 33.5 32.5
-AdaContrast (Chen et al. 2022)
-42.3 36.8 38.6 27.7 40.1 29.1 27.5 32.9 30.7 38.2 25.9 28.3 33.9 33.3 36.2 33.4
-Surgical
-fine-tuning (Lee et al. 2022)
-40.7 36.3 37.8 25.5 39.0 27.6 25.2 32.2 30.4 37.3 23.7 26.8 32.3 27.6 37.2 32.0
-DSS(Wang et al. 2024)
-39.7 36.0 37.2 26.3 35.6 27.5 25.1 31.4 30.0 37.8 24.2 26.0 30.0 26.3 31.1 30.9
-LAW (Park et al. 2024)
-41.0 36.7 38.3 25.6 37.0 27.8 25.2 30.7 30.0 37.3 24.4 27.6 31.2 27.8 34.9 31.7
-Ours
-37.3 32.5 34.9 26.2 35.3 27.5 24.6 28.8 29.1 34.1 23.5 26.9 31.2 26.6 34.1 30.1
+Source 73.0 68 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
+BN-1 42.1 40.7 42.7 27.6 41.9 29.7 27.9 34.9 35.0 41.5 26.5 30.3 35.7 32.9 41.2 35.4
+Tent Cont. 37.2 35.8 41.7 37.9 51.2 48.3 48.5 58.4 63.7 71.1 70.4 82.3 88.0 88.5 90.4 60.9
+CoTTA 40.1 37.7 39.7 26.9 38.0 27.9 26.4 32.8 31.8 40.3 24.7 26.9 32.5 28.3 33.5 32.5
+AdaContrast 42.3 36.8 38.6 27.7 40.1 29.1 27.5 32.9 30.7 38.2 25.9 28.3 33.9 33.3 36.2 33.4
+SFT (cont.) 40.7 36.3 37.8 25.5 39.0 27.6 25.2 32.2 30.4 37.3 23.7 26.8 32.3 27.6 37.2 32.0
+DSS 39.7 36.0 37.2 26.3 35.6 27.5 25.1 31.4 30.0 37.8 24.2 26.0 30.0 26.3 31.1 30.9
+LAW 41.0 36.7 38.3 25.6 37.0 27.8 25.2 30.7 30.0 37.3 24.4 27.6 31.2 27.8 34.9 31.7
+Ours 37.3 32.5 34.9 26.2 35.3 27.5 24.6 28.8 29.1 34.1 23.5 26.9 31.2 26.6 34.1 30.1
 Table 2: Mean errors (%) on CIFAR-100C - CTTA mean errors of the 15 corruptions (tasks) at a severity level of 5.
@@
 As the evaluation metric, we report the mean classification error.
-CTTA implementation details: To maintain fairness regarding the source models, we employ WideResNet-28 (Zagoruyko and Komodakis 2016) for CIFAR-10C, ResNeXt-29 (Xie et al.
-2017) for CIFAR-100C, and ResNet-50 (He et al.
+CTTA implementation details: To maintain fairness regarding the source models, we employ WideResNet-28 (36.4M params) (Zagoruyko and Komodakis 2016) for CIFAR-10C, ResNeXt-29 (6.8M params) (Xie et al.
+2017) for CIFAR-100C, and ResNet-50 (23.5M params) (He et al.
 2016) for ImageNet-C, all available on RobustBench (Croce et al.
@@
 Baselines: We compare against several CTTA baselines, including BN Stats Adapt (BN-1) (Schneider et al.
-2020), TENT Continual (Wang et al.
+2020), Tent Cont.
+(Wang et al.
 2020), CoTTA (Wang et al.
@@
 Following the findings of (Lee et al.
-2022), we continually fine-tune only the first convolutional block of the respective models, denoted as “Surgical fine-tuning (continual)”, using an SGD optimizer with fixed learning rates of 1e-3 for the CIFAR datasets and 5e-4 for ImageNet-C.
+2022), we continually fine-tune only the first convolutional block of the respective models, denoted as “SFT (cont.)”, using an SGD optimizer with fixed learning rates of 1e-3 for the CIFAR datasets and 5e-4 for ImageNet-C.
 For the results of LAW, we reproduce their results strictly following their official implementation  [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
@@
 Pixelate JPEG Mean
-Source
-97.8 97.1 98.2 81.7 89.8 85.2 78.0 83.5 77.1 75.9 41.3 94.5 82.5 79.3 68.6 82.0
-BN-1 (Schneider et al. 2020)
-85.0 83.7 85.0 84.7 84.3 73.7 61.2 66.0 68.2 52.1 34.9 82.7 55.9 51.3 59.8 68.6
-TENT Continual (Wang et al. 2020)
-81.6 74.6 72.7 77.6 73.8 65.5 55.3 61.6 63.0 51.7 38.2 72.1 50.8 47.4 53.3 62.6
-CoTTA (Wang et al. 2022)
-84.7 82.1 80.6 81.3 79.0 68.6 57.5 60.3 60.5 48.3 36.6 66.1 47.2 41.2 46.0 62.7
-EcoTTA (Song et al. 2023)
-- - - - - - - - - - - - - - - 63.4
-AdaContrast (Chen et al. 2022)
-82.9 80.9 78.4 81.4 78.7 72.9 64.0 63.5 64.5 53.5 38.4 66.7 54.6 49.4 53.0 65.5
-Surgical
-fine-tuning (Lee et al. 2022)
-83.4 79.5 77.2 86.1 82.4 72.2 61.1 63.1 64.7 51.2 35.9 72.9 56.2 46.7 50.2 65.5
-DSS (Wang et al. 2024)
-84.6 80.4 78.7 83.9 79.8 74.9 62.9 62.8 62.9 49.7 37.4 71.0 49.5 42.9 48.2 64.6
-LAW (Park et al. 2024)
-81.8 75.0 72.1 77.4 73.2 63.9 54.6 57.9 61.0 50.1 36.3 68.6 49.0 46.2 49.1 61.0
-Ours
-81.1 73.3 70.8 77.0 71.8 62.3 53.9 56.7 60.8 50.4 36.2 65.9 48.0 45.2 48.0 60.1
+Source 97.8 97.1 98.2 81.7 89.8 85.2 78.0 83.5 77.1 75.9 41.3 94.5 82.5 79.3 68.6 82.0
+BN-1 85.0 83.7 85.0 84.7 84.3 73.7 61.2 66.0 68.2 52.1 34.9 82.7 55.9 51.3 59.8 68.6
+Tent Cont. 81.6 74.6 72.7 77.6 73.8 65.5 55.3 61.6 63.0 51.7 38.2 72.1 50.8 47.4 53.3 62.6
+CoTTA 84.7 82.1 80.6 81.3 79.0 68.6 57.5 60.3 60.5 48.3 36.6 66.1 47.2 41.2 46.0 62.7
+EcoTTA - - - - - - - - - - - - - - - 63.4
+AdaContrast 82.9 80.9 78.4 81.4 78.7 72.9 64.0 63.5 64.5 53.5 38.4 66.7 54.6 49.4 53.0 65.5
+SFT (cont.) 83.4 79.5 77.2 86.1 82.4 72.2 61.1 63.1 64.7 51.2 35.9 72.9 56.2 46.7 50.2 65.5
+DSS 84.6 80.4 78.7 83.9 79.8 74.9 62.9 62.8 62.9 49.7 37.4 71.0 49.5 42.9 48.2 64.6
+LAW 81.8 75.0 72.1 77.4 73.2 63.9 54.6 57.9 61.0 50.1 36.3 68.6 49.0 46.2 49.1 61.0
+Ours 81.1 73.3 70.8 77.0 71.8 62.3 53.9 56.7 60.8 50.4 36.2 65.9 48.0 45.2 48.0 60.1
 Table 3: Mean errors (%) on ImageNet-C - CTTA mean errors of the 15 corruptions (tasks) at a severity level of 5.
@@
 Our method performs the best across all training setups and baselines for both datasets.
-We achieve large improvements of 4.9% and 5.2% for CIFAR-10C and 5.3% and 30.8% for CIFAR-100C, over BN-1 and TENT Continual respectively.
+We achieve large improvements of 4.9% and 5.2% for CIFAR-10C and 5.3% and 30.8% for CIFAR-100C, over BN-1 and Tent Cont.
+respectively.
 Concerning recent works of LAW and DSS, our method improves by 1.7% and 0.5% on CIFAR-10C and 1.6% and 0.8% on CIFAR-100C, respectively.
-TENT Continual excels initially but accumulates errors over time, increasing mean error.
+Tent Cont.
+excels initially but accumulates errors over time, increasing mean error.
 In CoTTA, the method’s reliance on full model updates renders it computationally expensive and unsuitable for deployment.
-An interesting observation is that with a drastic reduction in the number of trainable parameters, as in “Surgical fine-tuning (continual)”, the mean errors are better than most CTTA benchmarks, if not the best.
+An interesting observation is that with a drastic reduction in the number of trainable parameters, as in SFT (cont.), the mean errors are better than most CTTA benchmarks, if not the best.
 LAW uses the uncertain pseudo labels to calculate the domain shift approximated via the FIM.
@@
 1).
-This affirms the findings in (Lee et al.
+This confirms the findings in SFT (Lee et al.
 2022) that concentrated adaptation, for image corruptions, in the initial layers suffices.
 Results on Imagenet-C: In Table 3, we report the CTTA results on the ImageNet-C dataset.
-As shown, our method outperforms all the previous CTTA baselines and “Surgical fine-tuning (continual)”.
+As shown, our method outperforms all the previous CTTA baselines and SFT (cont.).
 We observe an improvement of 0.9% and 4.5% over LAW and DSS respectively.
-In addition, we achieve 2.6%, 5.4%, and 2.5% improvements over CoTTA, AdaContrast, and TENT Continual, respectively.
+In addition, we achieve 2.6%, 5.4%, and 2.5% improvements over CoTTA, AdaContrast, and Tent Cont., respectively.
       Dataset Source
 BN-1
-Adapt
 TENT
-Continual
+Cont.
 CoTTA AdaContrast
-Surgical
-fine-tuning
+SFT
 LAW Ours
@@
 ImageNet-C 58.4 48.3 46.4 38.8 66.3 41.4 40.5 38.5
-Table 4: Mean errors (in %) on CIFAR-10C, CIFAR-100C, and ImageNet-C - We report the GTTA mean errors averaged across all the severity levels (1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5).
+Table 4: Mean errors (in %) on CIFAR-10C, CIFAR-100C, and ImageNet-C - GTTA mean errors averaged across all severity levels (1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5).
 GTTA Classification Results
@@
 With a single NVIDIA A5000 GPU, PALM incurs a slightly higher adaptation time/batch, due to the two-stage proposed method.
-For example, on CIFAR-10, PALM runs for 0.8s per batch, whereas Surgical Fine-tuning (Surgical) and LAW take 0.45 s and 0.73 s respectively.
+For example, on CIFAR-10, PALM runs for 0.8s per batch, whereas SFT (cont.) and LAW take 0.45 s and 0.73 s respectively.
 Our method alleviates CF by only adapting a small subset of model parameters, ensuring that much of the prior source knowledge is preserved.
@@
 After adaptation, PALM retained source knowledge better, with mean accuracies of 92.29%, 77.46%, and 70.87%, compared to LAW’s 91.67%, 77.69%, and 71.57%.
-Despite accuracy decline with more tasks, especially after adaptation on ImageNet-C, PALM matches or outperforms LAW.
-Dataset Method
-The average % of
-adapted parameters/task
-(source parameters)
-CIFAR-10C Ours 4.34 (of 36.4M)
-CoTTA 100 (of 36.4M)
-CIFAR-100C Ours 6.65 (of 6.8M)
-CoTTA 100 (of 6.8M)
-ImageNet-C Ours 17.21 (of 23.5M)
-CoTTA 100 (of 23.5M)
-Table 5: Average % of adapted parameters/task.
-Adaptive LR optimizers vs Ours: Here, we study the effect of directly using adaptive LR optimizers like LARS (You, Gitman, and Ginsburg 2017), Prodigy (Mishchenko and Defazio 2023), and Adafactor (Shazeer and Stern 2018), for CTTA.
-We report the results in Table 6.
-The mentioned optimizers do not adjust layer-wise LRs based on parameter sensitivity, failing to capture rapid domain shifts, resulting in poorer performance compared to our method.
+Despite the accuracy decline with more tasks, especially after adaptation on ImageNet-C, PALM matches or outperforms LAW.
+      Dataset Ours CoTTA
+CIFAR-10C 4.34 100
+CIFAR-100C 6.65 100
+ImageNet-C 17.21 100
+Table 5: The average % of adapted source parameters/task.
       Dataset Ours LARS Prodigy Adafactor
@@
 Table 6: CTTA mean errors (%) using adaptive LR optimizers vs. ours.
+Adaptive LR optimizers vs Ours: Here, we study the effect of directly using adaptive LR optimizers like LARS (You, Gitman, and Ginsburg 2017), Prodigy (Mishchenko and Defazio 2023), and Adafactor (Shazeer and Stern 2018), for CTTA.
+We report the results in Table 6.
+The mentioned optimizers do not adjust layer-wise LRs based on parameter sensitivity, failing to capture rapid domain shifts and resulting in poorer performance compared to our method.
             WideResNet-28              ResNeXt-29              ResNet-50
@@
 Refer to caption Refer to caption Refer to caption
-Figure 2: Illustrations of LR importance $i_{j,n}^{t}$ across different convolutional blocks/stages/layers during CTTA.
+Figure 2: Illustrations of LR importance $i_{j,n}^{t}$ across different convolutional blocks(B)/stages(S)/layers(L) during CTTA.
 [Top row] - Variations for the domain “glass_blur” across datasets.
@@
 We achieve improvements of 2.4%, 3.7%, and 8.5% over CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
-Refer to caption Refer to caption
-a) Ablation on $\alpha$. b) Ablation on temperature T.
-Figure 3: Ablation results on smoothing factor $\alpha$ and temperature T.
       Dataset Pseudo-labels Ours
@@
 ImageNet-C 68.6 60.1
-Table 7: Layer importance using pseudo-labels vs ours.
+Table 7: Prediction uncertainty using pseudo-labels vs. ours.
              $\lambda$ 0 0.01 0.05 0.1 0.5 1.0
@@
 Table 8: Ablation results with varying $\lambda$.
-Contribution of smoothing factor $\alpha$: A key component of PALM is computing the moving average of sensitivities across data batches, smoothed by $\alpha$ (Eq.
-5).
-Figure 3(a) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$.
-We observe that the “domain-level” sensitivity across past batches of data is indeed useful.
-For $\alpha$ = 0 i.e., when $S_{j,n}^{t}$ of a selected parameter at timestep t is not considered, the respective datasets’ error rate drastically spikes.
-For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.7% over LAW.
-For CIFAR-100C and ImageNet-C, we see improvements of 1.6% and 0.9% over LAW, at $\alpha$ = 0.9 and 0.5 respectively.
-With higher values of $\alpha$, the results do not vary too much.
-Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{j,n}^{t}$) and current sensitivity ($S_{j,n}^{t}$) is needed.
-Effectiveness of temperature T: In our method, we compute the KL divergence between a uniform distribution u and a smoothed classifier’s output logits (Eq.
+Refer to caption Refer to caption
+a) Ablation on temperature T b) Ablation on $\alpha$
+Figure 3: Ablation results on smoothing factor $\alpha$ and temperature T.
+Effectiveness of temperature T: In our method, we compute the KL divergence between the uniform distribution u and a smoothed classifier’s output logits (Eq.
 1), denoted by $\hat{h}(x_{k})$.
-To study its influence, we illustrate the results in Figure 3(b).
+We illustrate the results in Figure 3(a) to study its influence.
 Ideally, larger values of T push the distributions of the predictions closer to u (Hinton, Vinyals, and Dean 2015) by spreading the probability among inputs.
@@
 For the CIFAR-C datasets, the errors begin to rise for T $\geq$ 100 which could be due to large spreads of the distributions.
-Robustness to regularization $\lambda$: Here, we understand the effect of the parameter $\lambda$ in Eq.
-11.
-We set $\lambda$in {0, 0.01, 0.05, 0.1, 0.5, 1.0}.
+Contribution of smoothing factor $\alpha$: A key component of PALM is computing the moving average of sensitivities across data batches, smoothed by $\alpha$ (Eq.
+5).
+Figure 3(b) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$.
+We observe that the “domain-level” sensitivity across past batches of data is indeed useful.
+For $\alpha$ = 0 i.e., when $S_{j,n}^{t}$ of a selected parameter at timestep t is not considered, the respective datasets’ error rate drastically spikes.
+For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.7% over LAW.
+For CIFAR-100C and ImageNet-C, we see improvements of 1.6% and 0.9% over LAW, at $\alpha$ = 0.9 and 0.5 respectively.
+With higher values of $\alpha$, the results do not vary too much.
+Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{j,n}^{t}$) and current sensitivity ($S_{j,n}^{t}$) is needed.
+Robustness to $\lambda$: In Eq.
+11, we set $\lambda$in {0, 0.01, 0.05, 0.1, 0.5, 1.0}.
 Table 8 reports the average classification errors on the benchmark datasets.
@@
 In addition, we observe that our method is very robust to $\lambda$ i.e., with subtle variations to differing $\lambda$.
-6 Conclusion
+Conclusion
 In this paper, we propose a novel approach called PALM, for continual test-time adaptation.
@@
 This effectively eliminates the need for pseudo-labels.
-For the parameters of these layers, selected based on the norm of the gradients, the learning rates are adjusted based on their sensitivity to domain shift.
+For the parameters of these layers, selected based on the norm of the gradients, the LRs are adjusted based on their sensitivity to domain shift.
 Extensive experiments are conducted on benchmark datasets to demonstrate the efficacy of our method against different baselines.
 The results reveal that PALM achieves superior performance and pushes the state-of-the-art on several standard baselines.
-Acknowledgments. We thank all the reviewers for their helpful comments and suggestions. This project was supported by a grant from the University of Texas at Dallas, Richardson, TX, USA.
+Acknowledgments
+We thank all the reviewers for their helpful comments and suggestions. This project was supported by a grant from the University of Texas at Dallas, Richardson, TX, USA.